### v2

# font=small

Abstract.
Weak Kleene logics are three-valued logics characterized by the presence of an infectious truth-value. In their external versions, as they were originally introduced by Bochvar [4] and Halldén [30], these systems are equipped with an additional connective capable of expressing whether a formula is classically true. In this paper we further expand the expressive power of external weak Kleen logics by modalizing them with a unary operator. The addition of an alethic modality gives rise to the two systems $\mathsf{B}_{\mathrm{e}}^{\square}$ and $\mathsf{PWK}^{\Box}_{\mathrm{e}}$, which have two different readings of the modal operator. We provide these logics with a complete and decidable Hilbert-style axiomatization w.r.t. a three-valued possible worlds semantics. The starting point of these calculi are new axiomatizations for the non-modal bases $\mathsf{B}_{\mathrm{e}}$ and $\mathsf{PWK}_{\mathrm{e}}$, which we provide using the recent algebraization results about these two logics. In particular, we prove the algebraizability of $\mathsf{PWK}_{\mathrm{e}}$. Finally some standard extensions of the basic modal systems are provided with their completeness results w.r.t. special classes of frames.
Key words and phrases:
Modal logic, Kleene logic, Bochvar external logic, PWK external logic.
2020 Mathematics Subject Classification:
Primary: 03B45. Secondary: 03B60.
1. Introduction
Modal logics are formalisms originally devised to speak about necessity and possibility of formulas belonging to classical propositional, or first-order, logic. They have been successfully extended to non-classical logics, namely modalities can be applied to formulas obeying rules of non-classical propositional (or first-order) logics. The process has involved a large variety of non-classical logics, including (but not limited to) intuitionistic logic [22],[23],[27], strong Kleene logic [24],[25], Belnap-Dunn [35], various fuzzy logics [29], [31], [3], [13], [40], and, more in general, the realm of substructural logics [11].
This tendency has only marginally involved weak Kleene logics. These (three-valued) logics, originally introduced by Bochvar [4], and subsequently investigated by Kleene [32], to deal with mathematical paradoxes and formulas possibly referring to non-existing objects and/or incorrectly written computer programs, have attracted much attentions in the recent past from several points of view: semantical [17], algebraic [7], [9], epistemic [38], [5], computer-theoretic [15], [18], topic-theoretic [1], and belief revision [14]. To the best of our knowledge, the only existing proposals of modal logics based on (some) weak Kleene logics are due to the works of Correia [19] and Segerberg [37]. Both papers focus on the modalized extension of Paraconsistent weak Kleene. [^1] Actually, Correia [19] introduces also the idea of a modalized version of Bochvar logic, but gives no axiomatization nor semantical analysis for that. While Correia’s work introduces an axiomatization and a relational semantics for the modal version of PWK, Segerberg’s [37] is based on the external version of Paraconsistent Weak Kleene and takes in consideration the existing difference between truth and non-falsity, within the three-valued realm, only with reference to establishing the semantical interpretation of modal formulas in a three-valued relational settings, while forgetting that such distinction is already made clear in the choice of the designated truth-values, leading either to Bochvar logic (whose consequence relation preserves truth only) or to Paraconsistent weak Kleene (or, Halldén logic, preserving non-falsity). Segerberg [37] essentially defines two different necessity operators on the external version of Paraconsistent weak Kleene. In the present work, inspired by Segerberg’s intuitions but guided by the above explained distinction, we will introduce modal logics, with a unique necessity operator, on both (the external versions of) Bochvar and Paraconsistent weak Kleene. Remarkably, the work of Segerberg relies on the external version of (a) weak Kleene logic. External weak Kleene logics are defined in an extended language from that of (propositional) classical logic, usually adopted also for weak Kleene, which we may now refer to as “internal Kleene logics”. The founding fathers of these formalisms – Bochvar and Halldén [4], [30] – actually defined their logics in the external language, which allows for a significant enrichment in expressiveness and the chance of recovering (propositional) classical logic, for a fragment of language, still working within a three-valued semantics. External Kleene logics turn out to be mathematically more interesting (than weak Kleene logics, as usually intended) as they are algebraizable in the sense of Blok and Pigozzi (see e.g. [26], Definition 3.11). This is part of our motivation to study modal Kleene logics in the external language: although the present work does not take in consideration the global version of modal logics over a certain relational structure, but only their local versions, we still expect to provide a modal basis on which algebraizability can be carried over from the propositional level (if one considers the global consequence relations instead of the local ones). Moreover, as weak Kleene logics turned out to be a particular case of a more general phenomenon, that of the “logics of variable inclusion” [9] – “internal” modal Kleene logics are indeed examples of logics of variable inclusion. As described in details in [9], these logics could be approached, syntactically, by imposing certain constraints on the inclusion of variables to standard modal logics and, semantically, by considering the construction of the Płonka sum of modal algebras or of its subvarieties “corresponding” to the extension of the modal logic $K$.
The main contribution of the present work is to introduce two different modal logics whose propositional bases are Bochvar external logic and the external version of Paraconsistent weak Kleene logic, respectively. For both of them, we provide Hilbert-style axiomatizations which are sound and complete with respect to a relational semantics, consisting of Kripke frames where the necessity operator is interpreted according to the choices of truth and non-falsity preservation imposed by the propositional basis of each logic. Finally, our work tries to shade a light on some extensions of weak Kleene modal logics, in particular, to those whose semantics is based on reflexive, transitive, and euclidean frames. The aim of this choice is that of providing useful tools for the analysis of epistemic concepts, such as knowledge, beliefs or ignorance, via non-classical logics (a tendency that has already be started e.g. in [6],[33]).
The paper is organized into six Sections (including this Introduction) plus and Appendix. In Section 2, we recall several important notions related to external Kleene logics that are needed to go through the paper. We also establish the algebraizability of PWK external logic, and use the algebraizability of both logics to provide new Hilbert-style axiomatizations, which we will argue have some advantages over the existing ones. In Section 3, we introduce, via a Hilbert-style axiomatization, a modal logic based on Bochvar external logic, for which we prove completeness and decidability with respect to the relational semantics. In section 4, we proceed analogously for PWK external logic. We dedicate Section 5 to introduce some axiomatic extensions of both modal Bochvar logic and modal PWK and show their completeness with respect to reflexive, transitive and euclidean relational models. Finally, in the Appendix (section 7), we prove the algebraizability of PWK external logic (with respect to the quasi-variety of Bochvar algebras as equivalent algebraic semantics).
2. External weak Kleene logics
Kleene three-valued logics are traditionally divided into two families, depending on the meaning given to the connectives: strong Kleene logics – counting strong Kleene [32] and the logic of paradox [34] – and weak Kleene logics, namely Bochvar logic and paraconsistent weak Kleene [7] (sometimes referred to as Halldén’s logic [30]). All the mentioned four logics are traditionally intended, and thus defined, over the (algebraic) language of classical logic. However, the intent of the first developer of the weak Kleene formalism, D. Bochvar, was to work within an enriched language allowing to express all classical “two-valued” formulas – which he referred to as external formulas – beside the genuinely “three-valued” ones. The result of this choice is the following language.
Definition 1.
Let us fix the language :$\mathcal{L}\colon\langle\neg,\lor,J_{{}_{2}},0,1\rangle$ of type $(1,2,1,0,0)$, which is obtained by enriching the classical language by an additional unary connective $J_{{}_{2}}$ (and the constants $0,1$), where the formula $J_{{}_{2}}\varphi$ is to be read as “$\varphi$ is true”. The language $\mathcal{L}$ will be referred to as external language, while its $J_{{}_{2}}$-reduct will be called internal. Let us refer to $\mathbf{Fm}$ as the formula algebra over $\mathcal{L}$, and to $Fm$ as its universe. More explicitly, formulas are inductively defined as follows:
$p\in Var$ |||| 00 |||| $1$ |||| $\neg\varphi$ |||| $\varphi\lor\psi$ |||| $J_{{}_{2}}\varphi.$
We will employ the following abbreviations: $\varphi\land\psi\coloneqq\neg(\neg\varphi\lor\neg\psi)$, $\varphi\to\psi\coloneqq\neg\varphi\lor\psi$, $\varphi\leftrightarrow\psi\coloneqq(\varphi\to\psi)\land(\psi\to\varphi)$, $J_{{}_{0}}\varphi\coloneqq J_{{}_{2}}\neg\varphi$, $J_{{}_{1}}\varphi\coloneqq\neg(J_{{}_{2}}\varphi\lor J_{{}_{2}}\neg\varphi)$, $+\varphi\coloneqq\neg J_{{}_{1}}\varphi$, and $\varphi\equiv\psi\coloneqq\displaystyle(J_{{}_{2}}\phi\leftrightarrow J_{{}_{2}}\psi)\land(J_{{}_{0}}\phi\leftrightarrow J_{{}_{0}}\psi)$.
The (algebraic) interpretation of the language $\mathcal{L}$ is given via the three-element algebra ${\mathbf{WK}}^{e}={\langle\{0,1,\nicefrac{{1}}{{2}}\},\neg,\lor,J_{{}_{2}},0,1\rangle}$ displayed in Figure 1.
$\lnot$
$1$ 00
/12 /12
00 $1$
$\lor$ 00 /12 $1$
00 00 /12 $1$
/12 /12 /12 /12
$1$ $1$ /12 $1$
$\varphi$ $J_{{}_{2}}\varphi$
$1$ $1$
/12 00
00 00
Figure 1. The algebra ${\mathbf{WK}}^{e}$.
The third value $\nicefrac{{1}}{{2}}$ is traditionally read as “meaningless” (see e.g. [20] and [39]) due to its infectious behavior. The defined connectives of conjunction and material implication have the expected tables:
$\land$ 00 /12 $1$
00 00 /12 00
/12 /12 /12 /12
$1$ 00 /12 $1$
$\to$ 00 /12 $1$
00 $1$ /12 $1$
/12 /12 /12 /12
$1$ 00 /12 $1$
Via $J_{{}_{2}}$ one can define other external connectives: $J_{{}_{0}}$ expressing the falseness of a formula, and $J_{{}_{1}}$ expressing its non-classicality. Moreover, $\equiv$ is the proper logical equivalence for the external language. Their interpretation in ${\mathbf{WK}}^{e}$ is displayed in the following tables. [^2] We are adopting here the notation due to Finn and Grigolia [21]; Bochvar [4] and Segerberg [36] instead used $\mathrm{t}$, $\mathsf{f}$, -- for $J_{{}_{2}}$, $J_{{}_{0}}$, $J_{{}_{1}}$, respectively.
$\varphi$ $J_{{}_{0}}\varphi$
$1$ 00
/12 00
00 $1$
$\varphi$ $J_{{}_{1}}\varphi$
$1$ 00
/12 $1$
00 00
$\equiv$ 00 /12 $1$
00 $1$ 00 00
/12 00 $1$ 00
$1$ 00 00 $1$
In the interpretation of the language $\mathcal{L}$ some formulas are evaluated into $\{0,1\}$ only (which is the universe of a Boolean subalgebra of ${\mathbf{WK}}^{e}$), by any homomorphism $h\colon\mathbf{Fm}\to{\mathbf{WK}}^{e}$: these are called external formulas (see Definition 4, for a syntactic definition).
2.1. Bochvar external logic
We recall that a logic $\mathsf{L}$ is induced by a matrix ${\langle\mathbf{M},F\rangle}$ (of the same language) when
$\Gamma\vdash_{\mathsf{L}}\varphi\text{ iff for all homomorphisms }h\colon\mathbf{Fm}\to\mathbf{M},$
$h[\Gamma]\subseteq F\text{ implies }h(\varphi)\in F.$
Definition 2.
Bochvar external logic $\mathsf{B}_{\mathrm{e}}$ is the logic induced by the matrix ${\langle{\mathbf{WK}}^{e},\{1\}\rangle}$. PWK external logic $\mathsf{PWK}_{\mathrm{e}}$ is the logic induced by the matrix ${\langle{\mathbf{WK}}^{e},\{1,\nicefrac{{1}}{{2}}\}\rangle}$.
Therefore, $\mathsf{B}_{\mathrm{e}}$ is the logic preserving only the value $1$ (for truth), while $\mathsf{PWK}_{\mathrm{e}}$ preserves both $1$ and $\nicefrac{{1}}{{2}}$ (thus, non-falsity). The latter, originally introduced by Halldén [30], has been later on studied by Segerberg [36], who named it $\mathrm{H_{0}}$. Both $\mathsf{B}_{\mathrm{e}}$ and $\mathsf{PWK}_{\mathrm{e}}$ are finitary logics, as they are defined by a finite set of finite matrices.
Hilbert-style axiomatizations of the two logics are given by Finn-Grigolia [21] and Segerberg [36], respectively. In order to introduce them, some technicalities are needed.
Definition 3.
An occurrence of a variable $x$ in a formula $\varphi$ is open if it does not fall under the scope of $J_{{}_{k}}$, for every $k\in\{0,1,2\}$. A variable $x$ in $\varphi$ is covered if all of its occurrences are not open, namely if for every occurrence of $x$ in $\varphi$ falls under the scope of $J_{{}_{k}}$, for some $k\in\{0,1,2\}$.
The intuition behind the notion of external formula is made precise by the following.
Definition 4.
A formula $\varphi\in Fm$ is called external if all its variables are covered.
The following axioms (and rule) define Finn and Grigolia’s Hilbert-style axiomatization [^3] To be precise, in [21] the following definition of the connective is given: $\varphi\equiv\psi\coloneqq\displaystyle\bigwedge_{i=0}^{2}J_{i}\varphi\leftrightarrow J_{{}_{i}}\psi$. Nonetheless, since the operator $J_{{}_{1}}$ is entirely determined by $J_{{}_{2}}$ and $J_{{}_{0}}$, Finn and Grigolia’s definition can be safely substituted with ours, obtaining an equivalent calculus (see e.g. [10]). of $\mathsf{B}_{\mathrm{e}}$ [21].
Axioms
  • (A1)
    $(\varphi\lor\varphi)\equiv\varphi$;
  • (A2)
    $(\varphi\lor\psi)\equiv(\psi\lor\varphi)$;
  • (A3)
    $((\varphi\lor\psi)\lor\chi)\equiv(\varphi\lor(\psi\lor\chi))$;
  • (A4)
    $(\varphi\land(\psi\lor\chi)\equiv((\varphi\land\psi)\lor(\varphi\land\chi))$;
  • (A5)
    ¬($\neg(\neg\varphi)\equiv\varphi$;
  • (A6)
    $\neg 1\equiv 0$;
  • (A7)
    $\neg(\varphi\lor\psi)\equiv(\neg\varphi\land\neg\psi)$;
  • (A8)
    $0\vee\varphi\equiv\varphi$;
  • (A9)
    $J_{{}_{2}}\alpha\equiv\alpha$;
  • (A10)
    $J_{{}_{0}}\alpha\equiv\neg\alpha$;
  • (A11)
    $J_{{}_{1}}\alpha\equiv 0$;
  • (A12)
    $J_{{}_{i}}\neg\varphi\equiv J_{{}_{2}-i}\varphi$, for any $i\in\{0,1,2\}$;
  • (A13)
    $J_{{}_{i}}\varphi\equiv\neg(J_{{}_{j}}\varphi\vee J_{{}_{k}}\varphi)$, with $i\neq j\neq k\neq i$;
  • (A14)
    $(J_{{}_{i}}\varphi\vee\neg J_{{}_{i}}\varphi)\equiv 1$, with $i\in\{0,1,2\}$;
  • (A15)
    $((J_{{}_{i}}\varphi\vee J_{{}_{k}}\psi)\land J_{{}_{i}}\varphi)\equiv J_{{}_{i}}\varphi$, with $i,k\in\{0,1,2\}$;
  • (A16)
    $(\varphi\lor J_{{}_{i}}\varphi)\equiv\varphi$, with $i\in\{1,2\}$;
  • (A17)
    $J_{{}_{0}}(\varphi\lor\psi)\equiv J_{{}_{0}}\varphi\land J_{{}_{0}}\psi$;
  • (A18)
    $J_{{}_{2}}(\varphi\lor\psi)\equiv(J_{{}_{2}}\varphi\land J_{{}_{2}}\psi)\vee(J_{{}_{2}}\varphi\land J_{{}_{2}}\neg\psi)\vee(J_{{}_{2}}\neg\varphi\land J_{{}_{2}}\psi)$.
Let $\alpha,\beta,\gamma$ denote external formulas only:
  • (A19)
    $\alpha\to(\beta\to\alpha)$;
  • (A20)
    $(\alpha\to(\beta\to\gamma))\to((\alpha\to\beta)\to(\alpha\to\gamma))$;
  • (A21)
    ($(\neg\alpha\to\neg\beta)\to(\beta\to\alpha)$.
Deductive rule
[…]
Notice that there is nothing special about the choice of axioms (A19)-(A21): it is only important that, together with modus ponens, yield a complete axiomatization for classical logic, but only relative to external formulas. [^4] Actually the original presentation in [21] contains a much longer axiomatization.
The fact that $\mathsf{B}_{\mathrm{e}}$ coincides with the logic induced by the above introduced Hilbert-style axiomatization has been proved in [6] (Finn and Grigolia [21, Theorem 3.4] only proved a weak completeness theorem for $\mathsf{B}_{\mathrm{e}}$). We will henceforth indicate by $\vdash_{\mathsf{B}_{\mathrm{e}}}$ both the consequence relation induced by the matrix ${\langle{\mathbf{WK}}^{e},\{1\}\rangle}$ and the one induced by the above Hilbert-style axiomatization.
The logic $\mathsf{B}_{\mathrm{e}}$ is algebraizable with the quasi-variety of Bochvar algebras as its equivalent algebraic semantics. [^5] This quasi-variety has been introduced by Finn and Grigolia [21], while its structural properties are studied in [10]. This means that there exists maps $\tau\colon Fm\to\mathcal{P}(Eq)$, $\rho\colon Eq\to\mathcal{P}(Fm)$ from formulas to sets of equations and from equations to sets of formulas such that
$\gamma_{1},\dots,\gamma_{n}\vdash_{\mathsf{B}_{e}}\varphi\iff\tau(\gamma_{1}),\dots,\tau(\gamma_{n})\vDash_{\mathcal{BCA}}\tau(\varphi)$
and
$\varphi\thickapprox\psi\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}\vDash_{\mathcal{BCA}}\tau\rho(\varphi\thickapprox\psi).$
The algebraizability of $\mathsf{B}_{\mathrm{e}}$ is witnessed by the transformers $\tau(\varphi)\coloneqq\{\varphi\thickapprox 1\}$ and $\rho(\varphi\thickapprox\psi)\coloneqq\{\varphi\equiv\psi\}$ (see [6] for details) and allows to provide a “standard” Hilbert-style axiomatization, whose axioms and rules make no difference bewteen external and non-external formulas. Recall that for a class $\mathcal{C}$ of algebras (of a certain type), the equational consequence relation $\Theta\vDash_{\mathcal{C}}\phi\approx\psi$ holds iff for all ${\mathbf{A}}\in\mathcal{C}$ and all homomorphisms (from the formulas in the same type) $h:\mathbf{Fm}\to{\mathbf{A}}$, if $h(\delta)=h(\epsilon)$ for all $\delta\approx\epsilon\in\Theta$, then $h(\phi)=h(\psi)$.
Using the equational description of the quasi-variety of Bochvar algebras presented in [10] (see in particular [10, Theorem 7]), we can apply the algorithm described in [26, Proposition 3.47], obtaining the following Hilbert-style calculus.
Definition 5.
A Hilbert-style axiomatization of $\mathsf{B}_{\mathrm{e}}$ is given by the following Axioms and Rules.
Axioms
  • ($\rho$-B1)
    $\varphi\lor\varphi\equiv\varphi$;
  • ($\rho$-B2)
    $\varphi\lor\psi\equiv\psi\lor\varphi$;
  • ($\rho$-B3)
    $(\varphi\lor\psi)\lor\chi\equiv\varphi\lor(\psi\lor\chi)$;
  • ($\rho$-B4)
    $\varphi\lor 0\equiv\varphi$;
  • ($\rho$-B5)
    $\neg\neg\varphi\equiv\varphi$;
  • ($\rho$-B6)
    $\neg(\varphi\lor\psi)\equiv\neg\varphi\land\neg\psi$;
  • ($\rho$-B7)
    $\neg 1\equiv 0$;
  • ($\rho$-B8)
    $\varphi\land(\psi\lor\chi)\equiv(\varphi\land\psi)\lor(\varphi\land\chi)$;
  • ($\rho$-B9)
    $J_{{}_{0}}J_{{}_{2}}\varphi\leftrightarrow\neg J_{{}_{2}}\varphi$;
  • ($\rho$-B10)
    $J_{{}_{2}}\varphi\leftrightarrow\neg(J_{{}_{0}}\varphi\lor J_{{}_{1}}\varphi)$;
  • ($\rho$-B11)
    $J_{{}_{2}}\varphi\lor\neg J_{{}_{2}}\varphi\leftrightarrow 1$;
  • ($\rho$-B12)
    $J_{{}_{2}}(\varphi\lor\psi)\leftrightarrow(J_{{}_{2}}\varphi\land J_{{}_{2}}\psi)\lor(J_{{}_{2}}\varphi\land J_{{}_{0}}\psi)\lor(J_{{}_{0}}\varphi\land J_{{}_{2}}\psi)$.
Deductive rules
  • ($\rho$-B13)
    $J_{{}_{2}}\varphi\leftrightarrow J_{{}_{2}}\psi,J_{{}_{0}}\varphi\leftrightarrow J_{{}_{0}}\psi\vdash\varphi\equiv\psi$;
  • (BAlg3)
    $\varphi\dashv\vdash J_{{}_{2}}\varphi\leftrightarrow 1$.
The axiomatization is equivalent to Finn and Grigolia’s calculus, moreover it consists of a proper set of axiom schemata. In fact Finn and Grigolia impose a syntactic restriction on axioms (A19)-(A21), as a result those are not schemata, instead each point represents a countable set of schemata. On the contrary, the above axiomatization makes no distinction between external and non-external formulas, hence it enjoys a proper closure under substitution.
The following results recaps some basic properties of $\mathsf{B}_{\mathrm{e}}$ which will be used in the following sections.
Lemma 6.
The following facts hold in $\mathsf{B}_{\mathrm{e}}$:
  1. (1)
    $\varphi,\varphi\to\psi\vdash\psi$;
  2. (2)
    $\vdash\alpha\leftrightarrow J_{{}_{2}}\alpha$, for $\alpha$ external formula;
  3. (3)
    $\varphi\mathrel{\dashv\mkern 1.5mu\vdash}J_{{}_{2}}\varphi$;
  4. (4)
    If $\alpha$ is an external formula and a classical theorem then $\vdash\alpha$;
  5. (5)
    $\neg\varphi\mathrel{\dashv\mkern 1.5mu\vdash}J_{{}_{0}}\varphi$;
  6. (6)
    $\vdash J_{{}_{0}}\varphi\to\neg J_{{}_{2}}\varphi$;
  7. (7)
    $\vdash J_{{}_{0}}\varphi\land J_{{}_{2}}\varphi\to 0$;
  8. (8)
    $\vdash\neg J_{{}_{2}}\varphi\to J_{{}_{1}}\varphi\lor J_{{}_{0}}\varphi$;
  9. (9)
    $\vdash J_{{}_{1}}\varphi\leftrightarrow J_{{}_{1}}\neg\varphi$;
  10. (10)
    $\vdash J_{{}_{1}}\alpha\to 0$ for $\alpha$ external formula;
  11. (11)
    $\neg J_{{}_{2}}\neg\varphi\vdash J_{{}_{2}}\varphi\lor J_{{}_{1}}\varphi$.
where $\varphi\leftrightarrow\psi$ is an abbrevation for $(\varphi\to\psi)\land(\psi\to\varphi)$.
Finally, let us recall that $\mathsf{B}_{\mathrm{e}}$ has a deduction theorem, in the following form.
Theorem 7 (Deduction Theorem for $\mathsf{B}_{\mathrm{e}}$).
It holds that $\Gamma\vdash_{\mathsf{B}_{\mathrm{e}}}\varphi$ iff there exist some formulas $\gamma_{1},\dots,\gamma_{n}\in\Gamma$ such that $\vdash_{\mathsf{B}_{\mathrm{e}}}J_{{}_{2}}\gamma_{1}\wedge\dots\wedge J_{{}_{2}}\gamma_{n}\to J_{{}_{2}}\varphi$.
2.2. External paraconsistent weak Kleene logic
The Hilbert-style axiomatization for $\mathsf{PWK}_{\mathrm{e}}$, introduced by Segerberg [36] is the following.
Axioms
  1. (A1)
    $(\varphi\lor\varphi)\to\varphi$;
  2. (A2)
    $\varphi\to(\varphi\lor\psi)$;
  3. (A3)
    $(\varphi\lor\psi)\to(\psi\lor\varphi)$;
  4. (A4)
    $(\varphi\to\psi)\to((\gamma\lor\varphi)\to(\gamma\lor\psi))$;
  5. (A5)
    $(\varphi\land\psi)\to\lnot(\lnot\varphi\lor\lnot\psi)$;
  6. (A6)
    ¬($\lnot(\lnot\varphi\lor\lnot\psi)\to(\varphi\land\psi)$;
  7. (A7)
    $\varphi\to 1$;
  8. (A8)
    $0\to\varphi$;
  9. (A9)
    $\varphi\to J_{{}_{2}}\varphi$;
  10. (A10)
    $J_{{}_{2}}\varphi\to\neg J_{{}_{0}}\varphi$;
  11. (A11)
    $J_{{}_{2}}(\varphi\land\psi)\leftrightarrow J_{{}_{2}}\varphi\land J_{{}_{2}}\psi$;
  12. (A12)
    $J_{{}_{2}}(\varphi\vee\psi)\leftrightarrow((J_{{}_{2}}\varphi\land J_{{}_{2}}\psi)\vee(J_{{}_{2}}\varphi\land J_{{}_{0}}\psi)\vee(J_{{}_{0}}\varphi\land J_{{}_{2}}\psi))$.
Deductive rule
[…]
As before, there is nothing special behind the choice of the axioms (A1) to (A8): one can simply choose any set of axioms which, together with the (usual) rule of Modus Ponens yields an axiomatization of (propositional) classical logic. Observe that the rule [RMP] consists of a linguistic restriction of the standard rule of Modus Ponens: a fact that shall not surprise, as a very similar restricted rule has been introduced for an axiomatization of $\mathsf{PWK}$ (in the language of classical logic) [7]. However, providing the same logic with a “standard” calculus presenting no linguistic restriction is preferable (for instance, for internal $\mathsf{PWK}$, such Hilbert-style axiomatizations can be found in [28], [8]). As in the case of Bochvar, also for $\mathsf{PWK}_{\mathrm{e}}$, we will indicate by $\vdash_{\mathsf{PWK}_{\mathrm{e}}}$ both the consequence relation induced by the matrix ${\langle{\mathbf{WK}}^{e},\{1,\nicefrac{{1}}{{2}}\}\rangle}$ and the one induced by the above Hilbert-style axiomatization.
Like $\mathsf{B}_{\mathrm{e}}$, also the logic $\mathsf{PWK}_{\mathrm{e}}$ is algebraizable with quasi-variety of Bochvar algebras as its equivalent algebraic semantic. The transformers that witness the algebraizability are $\tau(\varphi)\coloneqq\{\neg J_{{}_{0}}\varphi\thickapprox 1\}$ and $\rho(\varphi\thickapprox\psi)\coloneqq\{\varphi\equiv\psi\}$. We leave the proof of this result in the Appendix (see Section 7). Observe that, although it is not very common (see [26, pp.121-122]), the same class of algebras can play the role of equivalent algebraic semantics for different logics (clearly, the algebraizability is given by different transformers).
This algebraizability allows us to apply the algorithm of [26] already employed for $\mathsf{B}_{\mathrm{e}}$, obtaining a Hilbert-style axiomatization alternative to Segerberg’s.
Definition 8.
A Hilbert-style axiomatization of $\mathsf{PWK}_{\mathrm{e}}$ is given by the following Axioms and Rules.
Axioms
  • ($\rho$-B1)
    $\varphi\lor\varphi\equiv\varphi$;
  • ($\rho$-B2)
    $\varphi\lor\psi\equiv\psi\lor\varphi$;
  • ($\rho$-B3)
    $(\varphi\lor\psi)\lor\chi\equiv\varphi\lor(\psi\lor\chi)$;
  • ($\rho$-B4)
    $\varphi\lor 0\equiv\varphi$;
  • ($\rho$-B5)
    $\neg\neg\varphi\equiv\varphi$;
  • ($\rho$-B6)
    $\neg(\varphi\lor\psi)\equiv\neg\varphi\land\neg\psi$;
  • ($\rho$-B7)
    $\neg 1\equiv 0$;
  • ($\rho$-B8)
    $\varphi\land(\psi\lor\chi)\equiv(\varphi\land\psi)\lor(\varphi\land\chi)$;
  • ($\rho$-B9)
    $J_{{}_{0}}J_{{}_{2}}\varphi\leftrightarrow\neg J_{{}_{2}}\varphi$;
  • ($\rho$-B10)
    $J_{{}_{2}}\varphi\leftrightarrow\neg(J_{{}_{0}}\varphi\lor J_{{}_{1}}\varphi)$;
  • ($\rho$-B11)
    $J_{{}_{2}}\varphi\lor\neg J_{{}_{2}}\varphi\leftrightarrow 1$;
  • ($\rho$-B12)
    $J_{{}_{2}}(\varphi\lor\psi)\leftrightarrow(J_{{}_{2}}\varphi\land J_{{}_{2}}\psi)\lor(J_{{}_{2}}\varphi\land J_{{}_{0}}\psi)\lor(J_{{}_{0}}\varphi\land J_{{}_{2}}\psi)$.
Deductive rules
  • ($\rho$-B13)
    $J_{{}_{2}}\varphi\leftrightarrow J_{{}_{2}}\psi,J_{{}_{0}}\varphi\leftrightarrow J_{{}_{0}}\psi\vdash\varphi\equiv\psi$;
  • (PWKAlg3)
    $\varphi\dashv\vdash\neg J_{{}_{0}}\varphi\leftrightarrow 1$.
The only difference between the new axiomatizations proposed for $\mathsf{B}_{\mathrm{e}}$ and $\mathsf{PWK}_{\mathrm{e}}$ relies on the rules (PWKAlg3) and (BAlg3), which is expected since the two logics have the same equivalent algebraic semantics and differ only for the $\tau$ transformer.
Lemma 9.
The following facts hold for the logic $\mathsf{PWK}_{\mathrm{e}}$:
  1. (1)
    $\alpha,\alpha\to\beta\vdash\beta$, for every $\alpha,\beta$ external formulas;
  2. (2)
    $\varphi\mathrel{\dashv\mkern 1.5mu\vdash}\neg J_{{}_{0}}\varphi$;
  3. (3)
    every theorem of classical logic is a theorem of $\mathsf{PWK}_{\mathrm{e}}$;
  4. (4)
    $\vdash\neg J_{{}_{0}}0\to 0$;
  5. (5)
    $\vdash\neg\varphi\to J_{{}_{0}}\varphi$;
  6. (6)
    $\vdash\alpha\leftrightarrow J_{{}_{2}}\alpha$, for $\alpha$ external formula;
  7. (7)
    $\varphi\vdash J_{{}_{2}}\varphi\lor J_{{}_{1}}\varphi$;
  8. (8)
    $\vdash J_{{}_{2}}\neg\varphi\leftrightarrow J_{{}_{0}}\varphi$;
  9. (9)
    $\vdash J_{{}_{1}}\varphi\leftrightarrow J_{{}_{1}}\neg\varphi$;
  10. (10)
    $\vdash\neg J_{{}_{0}}1$;
  11. (11)
    $\vdash\varphi\to J_{{}_{2}}\varphi$;
  12. (12)
    $\vdash J_{{}_{2}}\varphi\to+\varphi$.
$\mathsf{PWK}_{\mathrm{e}}$ has a Deduction Theorem very similar to $\mathsf{B}_{\mathrm{e}}$, by just adapting the statement of Theorem 14 (from truth) to non-falsity, in the obvious way suggested by external connectives.
Theorem 10 (Deduction Theorem for $\mathsf{PWK}_{\mathrm{e}}$).
It holds that $\Gamma\vdash_{\mathsf{PWK}_{\mathrm{e}}}\varphi$ iff there exist some formulas $\gamma_{1},\dots,\gamma_{n}\in\Gamma$ such that $\vdash_{\mathsf{PWK}_{\mathrm{e}}}\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\to\neg J_{{}_{0}}\varphi$.
3. Modal Bochvar logic
From now on, by $\mathbf{Fm}$ we will denote the formula algebra constructed over a denumerable set of propositional variables $Var$ in the language :$\mathcal{L}:\neg,\vee,J_{{}_{2}},\square,0,1$ of type ${\langle 1,2,1,1,0,0\rangle}$. The connectives $\land$,$\to$ , → are defined as usual, while recall that $J_{{}_{0}}\varphi$ and $J_{{}_{1}}\varphi$ are abbreviations for $J_{{}_{2}}\neg\varphi$ and $\neg(J_{{}_{2}}\varphi\lor J_{{}_{2}}\neg\varphi)$, respectively. Let, moreover, $\diamond\varphi$ be an abbreviation for $\neg\square\neg\varphi$. Our aim with the above introduced language is to define a (local) modal logic whose propositional basis is Bochvar external logic and whose interpretation of formulas $\square\varphi$, in a relational semantics, is that $\square\varphi$ holds in a state when $\varphi$ holds (is equal to 1) in all related states.
We introduce the logic ${\langle\mathbf{Fm},\vdash_{\mathsf{B}_{\mathrm{e}}^{\square}}\rangle}$ as induced by the following Hilbert-style axiomatization.
Axioms
  • the axioms of $\mathsf{B}_{\mathrm{e}}$ in Definition 5;
  • (B1)
    $\square(J_{2}\varphi\to J_{2}\psi)\to(\square J_{2}\varphi\to\square J_{2}\psi)$;
  • (B2)
    $+\varphi\leftrightarrow+\square\varphi$;
  • (B3)
    $J_{{}_{2}}\square\varphi\to\square J_{{}_{2}}\varphi$;
  • (B4)
    $J_{{}_{0}}\square\varphi\to\neg\square J_{{}_{0}}\neg\varphi$.
Deductive rules
  • ($\rho$-B13) $J_{{}_{2}}\varphi\leftrightarrow J_{{}_{2}}\psi,J_{{}_{0}}\varphi\leftrightarrow J_{{}_{0}}\psi\vdash\varphi\equiv\psi$;
  • (BAlg3) $\varphi\dashv\vdash J_{{}_{2}}\varphi\leftrightarrow 1$;
  • (N): if $\vdash\varphi$ then $\vdash\square\varphi$.
Throughout this Section, for ease of notation, we will write $\vdash$ instead of $\vdash_{\mathsf{B}_{\mathrm{e}}^{\square}}$ and refer to the above introduced logic simply as $\mathsf{B}_{\mathrm{e}}^{\square}$. The axiom (B1) can be generalized to all external formulas, as follows.
Lemma 11.
For $\alpha,\beta$ external formulas, the following is a theorem of $\mathsf{B}_{\mathrm{e}}^{\square}$:
  • (BK)
    $\square(\alpha\to\beta)\to(\square\alpha\to\square\beta)$
Proof.
Consider arbitrary external formulas $\alpha,\beta$. Let us start by instantiating (B1) as $\vdash\square(J_{{}_{2}}\alpha\to J_{{}_{2}}\beta)\to(\square J_{{}_{2}}\alpha\to\square J_{{}_{2}}\beta)$. Recalling Lemma 6, for $\gamma$ external it holds $\vdash\gamma\leftrightarrow J_{{}_{2}}\gamma$, therefore we can substitute equivalent formulas and obtain $\vdash\square(\alpha\to\beta)\to(\square\alpha\to\square\beta)$. ∎
Remark 12.
The rule of Modus Ponens (MP) obviously holds for $\mathsf{B}_{\mathrm{e}}^{\Box}$ (see Lemma 6).
Lemma 13.
In the logic $\mathsf{B}_{\mathrm{e}}^{\square}$ the following facts hold:
  1. (1)
    $J_{{}_{2}}\square\varphi\vdash J_{{}_{2}}\square J_{{}_{2}}\varphi$;
  2. (2)
    $\square(J_{i}\varphi\land J_{k}\psi)\leftrightarrow\square J_{i}\varphi\land\square J_{k}\psi$ for every $i,k\in\{0,1,2\}$;
  3. (3)
    $\diamond(J_{i}\varphi\vee J_{k}\psi)\leftrightarrow\diamond J_{i}\varphi\vee\diamond J_{k}\psi$ for $i,k\in\{0,1,2\}$;
  4. (4)
    $\vdash J_{{}_{2}}\diamond\varphi\to\diamond J_{{}_{2}}\varphi\lor\diamond J_{{}_{1}}\varphi$.
Proof.
Since $\mathsf{B}_{\mathrm{e}}^{\square}$ contains all axioms from $\mathsf{B}_{\mathrm{e}}$, we will freely make use of theorems and rules holding in the latter logic. In particular, notice that classical logic can always be employed on external formulas (by Lemma 6 and the fact that MP is a rule of $\mathsf{B}_{\mathrm{e}}$).
  1. (1)
    By (B3) and (MP), $J_{{}_{2}}\square\varphi\vdash\square J_{{}_{2}}\varphi$; by Lemma 6 (and the transitivity of $\vdash$) we get $\square J_{{}_{2}}\varphi\vdash J_{{}_{2}}\square J_{{}_{2}}\varphi$.
  2. (2)
    By classical logic, $\vdash J_{i}\varphi\land J_{k}\psi\to J_{i}\varphi$, by (N) $\vdash\square(J_{i}\varphi\land J_{k}\psi\to J_{i}\varphi)$. Applying (BK) and (MP), $\vdash\square(J_{i}\varphi\land J_{k}\psi)\to\square J_{i}\varphi$. By the same reasoning, $\vdash\square(J_{i}\varphi\land J_{k}\psi)\to\square J_{k}\psi$ as well. We conclude $\vdash\square(J_{i}\varphi\land J_{k}\psi)\to\square J_{i}\varphi\land\square J_{k}\psi$. For the other direction: $\vdash J_{i}\varphi\to(J_{i}\psi\to J_{i}\varphi\land J_{k}\psi)$ by classical logic. By (N), $\vdash\square(J_{i}\varphi\to(J_{k}\psi\to J_{i}\varphi\land J_{k}\psi))$, and by (BK) and (MP) twice, we have $\vdash\square J_{i}\varphi\to(\square J_{k}\psi\to\square(J_{i}\varphi\land J_{k}\psi))$. Again by classical logic $\vdash\square J_{i}\varphi\land\square J_{k}\psi\to\square(J_{i}\varphi\land J_{k}\psi)$.
  3. (3)
    By classical logic, $\vdash\neg J_{i}\varphi\land\neg J_{k}\psi\to\neg J_{i}\varphi$, by (N) ($\vdash\square(\neg J_{i}\varphi\land\neg J_{k}\psi\to\neg J_{i}\varphi)$. Applying (BK) and (MP), ($\vdash\square(\neg J_{i}\varphi\land\neg J_{k}\psi)\to\square\neg J_{i}\varphi$. By contraposition and de Morgan, $\vdash\neg\square\neg J_{i}\varphi\to\neg\square\neg(J_{i}\varphi\lor J_{k}\psi)$, which is the definition of $\vdash\diamond J_{i}\varphi\to\diamond(J_{i}\varphi\lor J_{k}\psi)$. By the same reasoning, $\vdash\diamond J_{k}\psi\to\diamond(J_{i}\varphi\lor J_{k}\psi)$ as well. We conclude $\vdash\diamond J_{i}\varphi\lor\diamond J_{k}\psi\to\diamond(J_{i}\varphi\lor J_{k}\psi)$. For the other direction, $\vdash\neg J_{i}\varphi\land\neg J_{k}\psi\to\neg(J_{i}\varphi\lor J_{k}\psi)$ by classical logic. By (N), ($\vdash\square(\neg J_{i}\varphi\land\neg J_{k}\psi\to\neg(J_{i}\varphi\lor J_{k}\psi))$, and by (BK) and (MP) twice we have ($\vdash\square(\neg J_{i}\varphi\land\neg J_{k}\psi)\to\square\neg(J_{i}\varphi\lor J_{k}\psi)$. By point (2) we can distribute box, ($\vdash(\square\neg J_{i}\varphi\land\square\neg J_{k}\psi)\to\square\neg(J_{i}\varphi\lor J_{k}\psi)$. By classical logic, ¬($\vdash\neg(\neg\square\neg J_{i}\varphi\lor\neg\square\neg J_{k}\psi)\to\square\neg(J_{i}\varphi\lor J_{k}\psi)$. By contraposition we conclude $\vdash\diamond(J_{i}\varphi\lor J_{k}\psi)\to\diamond J_{i}\varphi\lor\diamond J_{k}\psi$.
  4. (4)
    By (B4) $\vdash J_{{}_{0}}\Box\neg\varphi\to\neg\Box J_{{}_{0}}\varphi$. For the linguistic abbreviations introduced, we have that the antecedent $J_{{}_{0}}\Box\neg\varphi=J_{{}_{2}}\neg\Box\neg\varphi=J_{{}_{2}}\diamond\varphi$; while the consequent $\neg\Box J_{{}_{0}}\varphi=\neg\Box\neg(J_{{}_{2}}\varphi\lor J_{{}_{1}}\varphi)=\diamond(J_{{}_{2}}\varphi\lor J_{{}_{1}}\varphi)=\diamond J_{{}_{2}}\varphi\lor\diamond J_{{}_{1}}\varphi$.
Theorem 14 (Deduction Theorem).
For the logic $\mathsf{B}_{\mathrm{e}}^{\square}$, it holds that $\Gamma\vdash\varphi$ iff there exist some formulas $\gamma_{1},\dots,\gamma_{n}\in\Gamma$ such that $\vdash J_{{}_{2}}\gamma_{1}\wedge\dots\wedge J_{{}_{2}}\gamma_{n}\to J_{{}_{2}}\varphi$.
Proof.
The right to left direction is obvious. The other direction is proved by induction on the length of the derivation of $\varphi$ from $\Gamma$. We just show the inductive case of the rule (N). Let $\varphi=\Box\psi$, for some $\psi\in Fm$, and $\Gamma\vdash\Box\psi$, and the last deduction rule applied is (N), hence it holds $\vdash\psi$. By the latter fact, we have $\vdash\Box\psi$, hence $\vdash J_{{}_{2}}\Box\psi$ (by Lemma 6). By induction hypothesis, there exists some formulas $\gamma_{1},\dots,\gamma_{n}\in\Gamma$ such that $\vdash J_{{}_{2}}\gamma_{1}\wedge\dots\wedge J_{{}_{2}}\gamma_{n}\to J_{{}_{2}}\psi$. Since the axioms (and rule) of classical logic hold for external formulas (Lemma 6), we have $\vdash J_{{}_{2}}\Box\psi\to(J_{{}_{2}}\gamma_{1}\wedge\dots\wedge J_{{}_{2}}\gamma_{n}\to J_{{}_{2}}\Box\psi)$, hence, by (MP), $\vdash J_{{}_{2}}\gamma_{1}\wedge\dots\wedge J_{{}_{2}}\gamma_{n}\to J_{{}_{2}}\Box\psi$. ∎
3.1. Semantics
The intended semantics of this modal logic consists of a relational (Kripke-style) structure where formulas, in each world, are evaluated into ${\mathbf{WK}}^{e}$ (this has been already implemented for instance in [6]). We introduce these structures according to the current terminology adopted in many-valued modal logics.
Definition 15.
A weak three-valued Kripke model $\mathcal{M}$ is a structure ${\langle W,R,v\rangle}$ such that:
  1. (1)
    $W$ is a non-empty set (of possible worlds);
  2. (2)
    $R$ is a binary relation over $W$ ($R\subseteq W\times W$);
  3. (3)
    $v$ is a map, called valuation, assigning to each world and each variable, an element in ${\mathbf{WK}}^{e}$ ($v\colon W\times\mathbf{Fm}\to{\mathbf{WK}}^{e}$).
Non-modal formulas will be interpreted as in $\mathsf{B}_{\mathrm{e}}$, i.e. we assume that $v$ is a homomorphism, in its second component, with respect to $\neg,\vee,J_{{}_{2}},1,0$. The reduct $\mathcal{F}={\langle W,R\rangle}$ of a model $\mathcal{M}$ is called frame.
Notation: for ordered pairs of related elements, we equivalently write $(w,s)\in R$ or $wRs$.
The semantical interpretation of the modality $\Box$ is what characterize a special family of weak three-valued Kripke models:
Definition 16.
A Bochvar-Kripke model is a weak three-valued Kripke model ${\langle W,R,v\rangle}$ such that $v$ evaluates formulas of the form $\Box\varphi$ according to:
  • (1)
    $v(w,\Box\varphi)=1$ iff $v(w,\varphi)\neq\nicefrac{{1}}{{2}}$ and $v(s,\varphi)=1$ for every $s\in W$ such that $wRs$.
  • (0)
    $v(w,\Box\varphi)=0$ iff $v(w,\varphi)\neq\nicefrac{{1}}{{2}}$ and there exists $s\in W$ such that $wRs$ and $v(s,\varphi)\neq 1$.
  • ($\nicefrac{{1}}{{2}}$)
    $v(w,\Box\varphi)=\nicefrac{{1}}{{2}}$ iff $v(w,\varphi)=\nicefrac{{1}}{{2}}$.
To simplify, within this Section by model we intend Bochvar-Kripke model.
Remark 17.
Observe that, in every model ${\langle W,R,v\rangle}$ with $w\in W$, it holds that $v(w,\diamond\varphi)=1$ iff $v(w,\varphi)\neq\nicefrac{{1}}{{2}}$ and there exists $s\in W$ such that $wRs$ and $v(s,\varphi)\neq 0$, while $v(w,\diamond\varphi)=0$ iff $v(w,\varphi)\neq\nicefrac{{1}}{{2}}$ and $v(s,\varphi)=0$ for every $s\in W$ such that $wRs$.
As usual in modal logic, one can opt to study the local or the global consequence relation related to a class of frame. In this paper, we will always deal with the former. Accordingly, we denote by $\models^{l}_{\mathsf{B}_{\mathrm{e}}^{\square}}$ the local modal Bochvar external logic on the class of all frames obtained by taking $\{1\}$ as designated value, that is:
Definition 18.
$\Gamma\vDash^{l}_{\mathsf{B}_{\mathrm{e}}^{\square}}\varphi$ iff for all models ${\langle W,R,v\rangle}$ and all $w\in W$, if $v(w,\gamma)=1,\forall\gamma\in\Gamma$, then $v(w,\varphi)=1$.
[…]
Definition 19.
A formula $\varphi$ is satisfied (valid) in a model ${\langle W,R,v\rangle}$ if $v(w,\varphi)=1$, for some (all) $w\in W$. A formula $\varphi$ is valid in a frame $\mathcal{F}$ (notation $\mathcal{F}\vDash\varphi$) if it is valid in ${\langle\mathcal{F},v\rangle}$, for all valuations $v$. A formula $\varphi$ is valid in a class of frames $\mathcal{K}$ (notation $\mathcal{K}\vDash\varphi$) if it is valid in every frame $\mathcal{F}\in\mathcal{K}$.
3.2. Completeness and decidability
Definition 20.
A set $\Gamma\subset Fm$ is consistent if $\Gamma\not\vdash\varphi$, for some $\varphi\in Fm$. It is inconsistent if it is not consistent.
Remark 21.
Equivalently, a set $\Gamma\subset Fm$ is consistent if there is no formula $\varphi\in Fm$, such that $\Gamma\vdash J_{{}_{2}}\varphi$ and $\Gamma\vdash\neg J_{{}_{2}}\varphi$. Observe that this is equivalent to say that $\Gamma\not\vdash 0$.
Definition 22.
A consistent set $\Gamma$ is maximally consistent (or complete) whenever $\Gamma\subset\Gamma^{\prime}$ implies that $\Gamma^{\prime}$ is inconsistent. Equivalently, $\Gamma$ is maximally consistent iff, for every $\varphi\in Fm$ exactly one of the following holds:
  • i)
    $\varphi\in\Gamma$;
  • ii)
    $\neg\varphi\in\Gamma$;
  • iii)
    $J_{{}_{1}}\varphi\in\Gamma$;
Definition 23.
A formula $\varphi$ is meaningful in a maximally consistent set $w$ if $x\in w$ or $\neg x\in w$, for every open variable $x\in\varphi$.
Observe that the definition of meaningful formulas implies, semantically, that such formulas are those evaluated, in a state, into the two-elements Boolean algebra $\mathbf{B}_{2}$ only. The definition of meaningful formula obviously apply to variables as well.
Lemma 24.
Let $w$ be a maximally consistent set of formulas, then:
  1. (1)
    if $\neg\varphi\not\in w$ and all the variables occurring in $\varphi$ are meaningful in $w$, then $\varphi\in w$;
  2. (2)
    if all the variables of $\varphi$ are covered, then $\varphi$ is meaningful in $w$;
  3. (3)
    if $\vdash\varphi$ then $\varphi\in w$.
Proof.
We just show (3) (as the other claims can be found also in [37, Lemma 4.6]). Suppose that $\vdash\varphi$ and, by contradiction, that either $\neg\varphi\in w$ or $J_{{}_{1}}\varphi\in w$. Let us assume that $\neg\varphi\in w$. From $\vdash\varphi$ it follows $\varphi\in w$. By Lemma 6 we have $w\vdash J_{{}_{2}}\varphi$ and $w\vdash J_{{}_{0}}\varphi$. Applying the same lemma, the latter yields $w\vdash\neg J_{{}_{2}}\varphi$, in contradiction with the assumption that $w$ is (maximally) consistent (see Remark 21). One can reason similarly for the case $J_{{}_{1}}\varphi\in w$. ∎
Lemma 25.
[37, Lemma 4.7] Let $w$ be a maximally consistent set of formulas, t.f.a.e.
  1. (1)
    $\varphi$ is meaningful in $w$;
  2. (2)
    either $\varphi\in w$ or $\neg\varphi\in w$;
  3. (3)
    $+\varphi\in w$;
  4. (4)
    $\Box\varphi$ is meaningful in $w$;
  5. (5)
    $\diamond\varphi$ is meaningful in $w$.
Lemma 26.
[37, Lemma 4.8] For every maximally consistent set $w$ the following hold:
  1. (1)
    If $\varphi\to\psi\in w$ and $\varphi\in w$ then $\psi\in w$;
  2. (2)
    $\varphi\land\psi\in w$ if and only if $\varphi,\psi\in w$;
  3. (3)
    $\varphi\lor\psi\in w$ if and only if $\varphi\in w$ or $\psi\in w$;
  4. (4)
    $\varphi\in w$ if and only if $J_{{}_{2}}\varphi\in w$;
  5. (5)
    $J_{{}_{2}}\varphi\in w$ if and only if $\neg J_{{}_{2}}\varphi\not\in w$.
Lemma 27.
Let $\Gamma$ be a consistent set of formulas. $\Gamma\cup\{\varphi\}$ is inconsistent if and only if $\Gamma\vdash\neg\varphi$ or $\Gamma\vdash J_{{}_{1}}\varphi$.
Proof.
Let $\Gamma$ be a consistent set of formulas.
($\Rightarrow$) Let $\Gamma\cup\{\varphi\}$ be inconsistent and $\Gamma\not\vdash\neg\varphi$. By assumption, $\Gamma\cup\{\varphi\}\vdash 0$. By Theorem 14, there exist formulas $\gamma_{1},\dots,\gamma_{n}\in\Sigma$ such that $\vdash J_{{}_{2}}\gamma_{1}\wedge\dots\wedge J_{{}_{2}}\gamma_{n}\wedge J_{{}_{2}}\varphi\to J_{{}_{2}}0$, hence $\vdash J_{{}_{2}}\gamma_{1}\wedge\dots\wedge J_{{}_{2}}\gamma_{n}\to\neg J_{{}_{2}}\varphi$, thus $\Gamma\vdash\neg J_{{}_{2}}\varphi$. Since $\vdash\neg J_{{}_{2}}\varphi\to J_{{}_{1}}\varphi\lor J_{{}_{0}}\varphi$ by Lemma 6, $\Gamma\vdash J_{{}_{1}}\varphi\lor J_{{}_{0}}\varphi$. By assumption, $\Gamma\not\vdash\neg\varphi$ which implies $\Gamma\not\vdash J_{{}_{0}}\varphi$, hence $\Gamma\vdash J_{{}_{1}}\varphi$.
($\Leftarrow$) Let $\Gamma\vdash\neg\varphi$ or $\Gamma\vdash J_{{}_{1}}\varphi$. Suppose $\Gamma\vdash\neg\varphi$ is the case, hence $\Gamma\vdash J_{{}_{0}}\varphi$ (by Lemma 6). On the other hand, $\Gamma\cup\{\varphi\}\vdash J_{{}_{2}}\varphi$, hence $\Gamma\cup\{\varphi\}\vdash J_{{}_{0}}\varphi\land J_{{}_{2}}\varphi$, and since by Lemma 6 it is a theorem that $\vdash J_{{}_{0}}\varphi\land J_{{}_{2}}\varphi\to 0$, then $\Gamma\cup\{\varphi\}$ is inconsistent. The proof is analogue in case $\Gamma\vdash J_{{}_{1}}\varphi$. ∎
Lemma 28 (Lindenbaum’s Lemma).
Let $\Gamma$ be a consistent set of formulas such that $\Gamma\not\vdash\varphi$, for some $\varphi\in Fm$, then there exists a maximally consistent set of formulas $w$ such that $\Gamma\subseteq w$ and such that $\varphi\not\in w$.
Proof.
Consider an enumeration $\psi_{1},\psi_{2},\psi_{3},\dots$ of the formulas in $Fm$. Define:
$\Gamma_{0}=\begin{cases}\Gamma\cup\{\neg\varphi\}\text{ if consistent,}\\ \Gamma\cup\{J_{{}_{1}}\varphi\}\text{ otherwise.}\end{cases}$
$\Gamma_{i+1}=\begin{cases}\Gamma_{i}\cup\{\psi_{i}\}\text{ if consistent, else}\\ \Gamma_{i}\cup\{\neg\psi_{i}\}\text{ if consistent, else}\\ \Gamma_{i}\cup\{J_{{}_{1}}\psi_{i}\}.\end{cases}$
$w=\bigcup_{i\in\mathbb{N}}\Gamma_{i}.$
Observe that, by construction, $w$ is maximal. We want to show that $w$ is also consistent. We first claim that $\Gamma_{0}$ is consistent. If $\Gamma_{0}=\Gamma\cup\{\neg\varphi\}$ then it is consistent by construction. Differently, $\Gamma_{0}=\Gamma\cup\{J_{{}_{1}}\varphi\}$, which means that $\Gamma\cup\{\neg\varphi\}$ is inconsistent. Hence, by Lemma 27, $\Gamma\vdash\neg\neg\varphi$ or $\Gamma\vdash J_{{}_{1}}\neg\varphi$. However, $\Gamma\not\vdash\neg\neg\varphi$ (since, by assumption, $\Gamma\not\vdash\varphi$ and $\vdash_{\mathsf{B}_{\mathrm{e}}}\varphi\leftrightarrow\neg\neg\varphi$), so $\Gamma\vdash J_{{}_{1}}\neg\varphi$, which implies $\Gamma\vdash J_{{}_{1}}\varphi$ (as $\vdash_{\mathsf{B}_{\mathrm{e}}}J_{{}_{1}}\varphi\leftrightarrow J_{{}_{1}}\neg\varphi$). By Lemma 27, $\Gamma_{0}=\Gamma\cup\{J_{{}_{1}}\varphi\}$ is consistent if and only if $\Gamma\not\vdash\neg J_{{}_{1}}\varphi$ and $\Gamma\not\vdash J_{{}_{1}}J_{{}_{1}}\varphi$. Now, since $\Gamma$ is consistent and $\Gamma\vdash J_{{}_{1}}\varphi$, then $\Gamma\not\vdash\neg J_{{}_{1}}\varphi$. Moreover, since $\Gamma$ is consistent $\Gamma\not\vdash J_{{}_{1}}J_{{}_{1}}\varphi$ (as $\vdash J_{{}_{1}}J_{{}_{1}}\varphi\to 0$). This shows that $\Gamma_{0}$ is consistent.
We claim that $\Gamma_{i+1}$ is consistent, given that $\Gamma_{i}$ is. So, suppose that $\Gamma_{i}\cup\{\varphi\}$ and $\Gamma_{i}\cup\{\neg\varphi\}$ are inconsistent. Then, by Lemma 27, $\Gamma_{i}\vdash\neg\varphi$ or $\Gamma_{i}\vdash J_{{}_{1}}\varphi$, and, $\Gamma_{i}\vdash\neg\neg\varphi$ or $\Gamma_{i}\vdash J_{{}_{1}}\neg\varphi$. By consistency of $\Gamma_{i}$, the only possible case is that $\Gamma_{i}\vdash J_{{}_{1}}\varphi$ and $\Gamma_{i}\vdash J_{{}_{1}}\neg\varphi$, from which follows the consistency of $\Gamma\cup\{J_{{}_{1}}\varphi\}$ (indeed, if it is not consistent then, by Lemma 27, $\Gamma_{i}\vdash\neg J_{{}_{1}}\varphi$, in contradiction with the consistency of $\Gamma_{i}$). This shows that $w$ is maximal and consistent and, by construction, $\neg\varphi\in w$ or $J_{{}_{1}}\varphi\in w$, therefore $\varphi\not\in w$. ∎
As a first step to introduce canonical models, let us define the canonical relation.
Definition 29.
Let $\mathcal{W}$ be the set of all maximally consistent set of formulas. Then the canonical relation $\mathcal{R}\subset\mathcal{W}\times\mathcal{W}$ for $\mathsf{B}_{\mathrm{e}}^{\square}$ is defined, for every $w,s\in\mathcal{W}$ as:
$w\mathcal{R}s\text{ iff }\forall\varphi\in Fm\text{ s.t. }\Box\varphi\in w\text{ then }\varphi\in s.$
Lemma 30 (Existence Lemma).
For every maximally consistent set of formulas $w\in\mathcal{W}$ such that $\diamond\varphi\in w$ (for some $\varphi\in Fm$) then $\varphi$ is meaningful in $w$ and there exists a maximally consistent set $s\in\mathcal{W}$ such that $w\mathcal{R}s$ and either $\varphi\in s$ or $J_{{}_{1}}\varphi\in s$.
Proof.
Suppose that $\diamond\varphi\in w$, for some $w$ maximally consistent set of formulas. Consider the set
$s^{-}=\{J_{{}_{2}}\psi|\Box J_{{}_{2}}\psi\in w\}.$
Observe that $s^{-}\neq\emptyset$, as for every formula $\psi$ such that $\vdash\psi$, then $\vdash J_{{}_{2}}\psi$, hence $\vdash\Box J_{{}_{2}}\psi$, which implies $\Box J_{{}_{2}}\psi\in w$, by Lemma 24-(3). Let us show that $s^{-}$ is consistent. Suppose, by contradiction, that $s^{-}$ is inconsistent, then $s^{-}\vdash\gamma$, for every $\gamma\in Fm$, thus, in particular, $s^{-}\vdash\neg\varphi$. By Deduction Theorem there are formulas $J_{{}_{2}}\psi_{1},\dots,J_{{}_{2}}\psi_{n}\in s^{-}$ such that $\vdash J_{{}_{2}}J_{{}_{2}}\psi_{1}\land\dots\land J_{{}_{2}}J_{{}_{2}}\psi_{n}\to J_{{}_{2}}\neg\varphi$. Recall that $\vdash_{\mathsf{B}_{\mathrm{e}}}J_{{}_{2}}J_{{}_{2}}\gamma\leftrightarrow J_{{}_{2}}\gamma$, for every $\gamma\in Fm$, thus $\vdash J_{{}_{2}}\psi_{1}\land\dots\land J_{{}_{2}}\psi_{n}\to J_{{}_{2}}\neg\varphi$. By applying (N), we get $\vdash\Box(J_{{}_{2}}\psi_{1}\land\dots\land J_{{}_{2}}\psi_{n}\to J_{{}_{2}}\neg\varphi)$ and by distributing box ((BK) and Lemma 13), $\vdash\Box J_{{}_{2}}\psi_{1}\land\dots\land\Box J_{{}_{2}}\psi_{n}\to\Box J_{{}_{2}}\neg\varphi)$. Observe that, by construction of $s^{-}$, $\Box J_{{}_{2}}\psi_{i}\in w$, for every $i\in\{1,\dots,n\}$, hence $\Box J_{{}_{2}}\neg\varphi\in w$, i.e. $\neg\diamond\neg J_{{}_{2}}\neg\varphi\in w$. By Lemma 6, $\vdash\neg J_{{}_{2}}\neg\varphi\leftrightarrow J_{{}_{2}}\varphi\lor J_{{}_{1}}\varphi$, thus $\neg\diamond(J_{{}_{2}}\varphi\lor J_{{}_{1}}\varphi)\in w$, which implies (by distributivity of diamond, Lemma 13), $\neg(\diamond J_{{}_{2}}\varphi\lor\diamond J_{{}_{1}}\varphi)\in w$. On the other hand, $\diamond\varphi\in w$, hence $J_{{}_{2}}\diamond\varphi\in w$, which implies $\diamond J_{{}_{2}}\varphi\lor\diamond J_{{}_{1}}\varphi\in w$, by Lemma 13, giving raise to a contradiction with the fact that $w$ is consistent. Observe that we have also proved that $s^{-}\not\vdash\neg\varphi$, hence by Lindenbaum Lemma there exists a maximally consistent set $s$ such that $s^{-}\subseteq s$ and $\neg\varphi\not\in s$. By maximality, we have that either $\varphi\in s$ or $J_{{}_{1}}\varphi\in s$. To show that $w\mathcal{R}s$, suppose $\Box\gamma\in w$, for some $\gamma\in Fm$, then by Lemma 26 $J_{{}_{2}}\Box\gamma\in w$, hence, by (M3), $\Box J_{{}_{2}}\gamma\in w$, and by construction $J_{{}_{2}}\gamma\in s^{-}\subseteq s$, thus $\gamma\in s$ (by Lemma 26), showing that $w\mathcal{R}s$. Finally, let us show that $\varphi$ is meaningful in $w$. Since $\varphi\in w$, then $J_{{}_{2}}\varphi\in w$ and $+\varphi\in w$ (as $\vdash_{\mathsf{B}_{\mathrm{e}}}J_{{}_{2}}\varphi\to+\varphi$), namely that $\diamond\varphi$ is meaningful in $w$ and so is $\varphi$ (Lemma 25). ∎
We are ready to define the concept of canonical model.
Definition 31.
The canonical model for $\mathsf{B}_{\mathrm{e}}^{\square}$ is a model $\mathcal{M}={\langle\mathcal{W},\mathcal{R},v\rangle}$ where $\mathcal{W}$ is the set of all maximally consistent sets of formulas, $\mathcal{R}$ is the canonical relation for $\mathsf{B}_{\mathrm{e}}^{\square}$ and $v$ is defined as follows:
  • $v(w,x)=1$ if and only if $x\in w$;
  • $v(w,x)=0$ if and only if $\neg x\in X$;
  • $v(w,x)=\nicefrac{{1}}{{2}}$ if and only if $J_{{}_{1}}x\in w$,
for every $w\in\mathcal{W}$ and propositional variable $x$.
Lemma 32 (Truth Lemma).
Let $\mathcal{M}={\langle\mathcal{W},\mathcal{R},v\rangle}$ be the canonical model. Then, for every formula $\varphi\in Fm$ and every $w\in\mathcal{W}$, the following hold:
  1. (1)
    $v(w,\varphi)=1$ if and only if $\varphi\in w$;
  2. (2)
    $v(w,\varphi)=0$ if and only if $\neg\varphi\in w$;
  3. (3)
    $v(w,\varphi)=\nicefrac{{1}}{{2}}$ if and only if $J_{{}_{1}}\varphi\in w$.
Proof.
By induction on the length of the formula $\varphi$. We just show (1) for the inductive step when $\varphi=\Box\psi$, for some $\psi\in Fm$.
Observe that $v(w,\Box\psi)=1$ iff $\psi$ is meaningful in $w$ (i.e. $v(w,\psi)\neq\nicefrac{{1}}{{2}}$) and $\forall s$ s.t. $w\mathcal{R}s$, $v(s,\psi)=1$, thus, by induction hypothesis, iff $J_{{}_{1}}\psi\not\in w$ and $\psi\in s$ $\forall s$ s.t. $w\mathcal{R}s$.
()($\Rightarrow$)( ⇒ ) Suppose, by contradiction, that $v(w,\Box\psi)=1$ but $\Box\psi\not\in w$, hence, by maximality of $w$, $J_{{}_{1}}\Box\psi\in w$ or $\neg\Box\psi\in w$. Since $\psi$ is meaningful in $w$, so is $\Box\psi$ (Lemma 25), thus $J_{{}_{1}}\Box\psi\not\in w$. So, $\neg\Box\psi\in w$, i.e. $\diamond\neg\psi\in w$, hence, by Existence Lemma 30, there exists $s^{\prime}\in\mathcal{W}$ such that $w\mathcal{R}s^{\prime}$ such that $\neg\psi\in s^{\prime}$ or $J_{{}_{1}}\psi\in s^{\prime}$, which implies, by induction hypothesis, that $v(s^{\prime},\psi)=0$ or $v(s^{\prime},\psi)=\nicefrac{{1}}{{2}}$, a contradiction.
()($\Leftarrow$)( ⇐ ) Let $\Box\psi\in w$. Then $J_{{}_{2}}\Box\psi\in w$ (by Lemma 26) and $+\Box\psi\in w$ (since $\vdash_{\mathsf{B}_{\mathrm{e}}}J_{{}_{2}}\gamma\to+\gamma$). This means that $\Box\psi$ is meaningful in $w$, hence so is $\psi$. Moreover, for every $s\in\mathcal{W}$ such that $w\mathcal{R}s$, we have that $\psi\in s$ (by definition of $\mathcal{R}$), hence, by induction hypothesis, $v(s,\psi)=1$, from which $v(w,\psi)=1$. ∎
Theorem 33 (Completeness).
$\Gamma\vdash_{\mathsf{B}_{\mathrm{e}}}\varphi$ if and only if $\Gamma\models^{l}_{\mathsf{B}_{\mathrm{e}}}\varphi$.
Proof.
()($\Rightarrow$)( ⇒ ) It is easily checked that all the axioms are sound and the rules preserve soundness.
()($\Leftarrow$)( ⇐ ) Suppose $\Gamma\not\vdash\varphi$. Then $\Gamma$ is a consistent set of formulas, therefore, by the Lindenbaum Lemma 28, there exist a maximally consistent set $s$ such that $\Gamma\subseteq s$ and $\varphi\not\in s$, hence, by Truth Lemma 32, there exists a canonical countermodel, namely $(\mathcal{W},\mathcal{R},v)$, with $v(s,\gamma)=1$ for all $\gamma\in\Gamma$ and $v(s,\varphi)\neq 1$. ∎
In order to prove decidability for $\mathsf{B}_{\mathrm{e}}^{\square}$, we employ the filtration technique (see [2, pp. 77-80]). First we need to provide an extended notion of closure under subformulas.
Definition 34.
A set of formulas $\Sigma$ is closed under subformulas if $\forall\varphi,\psi\in\Sigma$:
  1. (1)
    if $\varphi\circ\psi\in\Sigma$ for any binary connective $\circ$, then $\varphi,\psi\in\Sigma$;
  2. (2)
    if $\neg\varphi\in\Sigma$ or $J_{{}_{2}}\varphi\in\Sigma$, then $\varphi\in\Sigma$;
  3. (3)
    if $\square\varphi\in\Sigma$, then $\varphi\in\Sigma$ and $+\varphi\in\Sigma$;
Notice that if a set of formulas is finite, its closure under subformulas is still finite.
Definition 35.
Let ${\langle W,R,v\rangle}$ be a model and $\Sigma$ be a finite set of formulas closed under subformulas. This set induces an equivalence relation over $W$ defined as follows: $w\equiv_{\Sigma}s$ iff $\forall\varphi\in\Sigma(v(w,\varphi)=1$ iff $v(s,\varphi)=1)$.
When the reference set $\Sigma$ is clear from the context, we denote the equivalence class $[w]_{\equiv_{\Sigma}}$ simply by $[w]$.
Definition 36.
Let $M={\langle W,R,v\rangle}$ be a model and $\Sigma$ be a finite set of formulas closed under subformulas. The filtration of $M$ through $\Sigma$ is the model ${\langle W^{f},R^{f},v^{f}\rangle}$ defined as:
  1. (1)
    $W^{f}=W/$$\equiv_{\Sigma}$;
  2. (2)
    $[w]R^{f}[s]$ iff $\exists w^{\prime}\in[w],s^{\prime}\in[s]$ s.t. $w^{\prime}Rs^{\prime}$;
  3. (3)
    $v^{f}([w],p)=1$ iff $v(w,p)=1$, for all variables $p\in\Sigma$.
Lemma 37.
Let ${\langle W^{f},R^{f},v^{f}\rangle}$ be a filtration of $M={\langle W,R,v\rangle}$ through $\Sigma$. For all $\varphi\in\Sigma,w\in W$, it holds $v(w,\varphi)=1$ iff $v^{f}([w],\varphi)=1$.
Proof.
By induction on the complexity of $\varphi\in\Sigma$. The Boolean cases are straightforward. Let $\varphi=J_{{}_{2}}\psi$, for some $\psi\in Fm$. $v(w,J_{{}_{2}}\psi)=1$ iff $v(w,\psi)=1$ iff, by induction hypothesis, $v^{f}([w],\psi)=1$ iff $v^{f}([w],J_{{}_{2}}\psi)=1$. Notice that by closure $\psi\in\Sigma$.
Let $\varphi=\square\psi$, for some $\psi\in Fm$. Suppose $v(w,\square\psi)=1$, which means that $v(w,+\psi)=1$ and for all $s\in W$ s.t. $wRs,v(s,\psi)=1$. Now $+\psi:=J_{{}_{2}}\psi\lor J_{{}_{2}}\neg\psi$, by the Boolean cases and the previous one we conclude $v^{f}([w],+\psi)=1$. By definition of filtration, $[w]R^{f}[s]$, and by induction hypothesis $v^{f}([s],\psi)=1$. Since this covers all the successors of $[w]$, then $v^{f}([w],\square\psi)=1$. Notice that by closure of $\psi$ under subformula, $J_{{}_{2}}\psi\lor J_{{}_{2}}\neg\psi\in\Sigma$. The other direction follows similarly. ∎
Theorem 38.
If a formula $\varphi$ is satisfiable in a model, it is satisfiable in a finite model.
Proof.
Let $\varphi$ be satisfied by a model $M={\langle W,R,v\rangle}$, and let $\Sigma$ be the closure under subformulas of $\{\varphi\}$. $\Sigma$ is finite. Now consider the filtration $M^{f}_{\Sigma}={\langle W^{f},R^{f},v^{f}\rangle}$ of $M$ through $\Sigma$. By theorem 37, $M^{f}_{\Sigma}$ satisfies $\varphi$. Consider the mapping $g:W^{f}\to\mathscr{P}(\Sigma)$ s.t. $g([w])=\{\psi|v(w,\psi)=1\}$. By definition of $\equiv_{\Sigma}$, $g$ is well-defined and injective. Denoting by $card(X)$ the cardinality of a set $X$, we have $card(W^{f})\leq card(\mathscr{P}(\Sigma))=2^{card(\Sigma)}$. ∎
Corollary 39 (Decidability).
The logic $\mathsf{B}_{\mathrm{e}}^{\Box}$ is decidable.
4. Modal $\mathsf{PWK}_{\mathrm{e}}$ logic
The modal extension of the propositional logic $\mathsf{PWK}_{\mathrm{e}}$ is defined over the same formula algebra $\mathbf{Fm}$ of $\mathsf{B}_{\mathrm{e}}^{\Box}$. The substantial (semantical) difference between $\mathsf{PWK}_{\mathrm{e}}^{\Box}$ and $\mathsf{B}_{\mathrm{e}}^{\Box}$ concern the interpretation of the modal formulas $\Box\varphi$, which follows the choice of the different truth-set in $\mathsf{PWK}_{\mathrm{e}}$: namely a modal formula $\Box\varphi$ will hold in a state $w$ iff it will also hold in all the related states $s$, namely in those the formula is not false.
The logic ${\langle\mathbf{Fm},\vdash_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}\rangle}$ is the consequence relation induced by the following Hilbert-style axiomatization.
Axioms
  • the axioms for $\mathsf{PWK}_{\mathrm{e}}$ introduced in Definition 8;
  • (P1)
    $\square(J_{2}\varphi\to J_{2}\psi)\to(\square J_{2}\varphi\to\square J_{2}\psi)$;
  • (P2)
    $\Box\varphi\leftrightarrow\Box\neg J_{{}_{0}}\varphi$;
  • (P3)
    $+\varphi\leftrightarrow+\square\varphi$.
Deductive rules
  • ($\rho$-B13) $J_{{}_{2}}\varphi\leftrightarrow J_{{}_{2}}\psi,J_{{}_{0}}\varphi\leftrightarrow J_{{}_{0}}\psi\vdash\varphi\equiv\psi$;
  • (PWKAlg3) $\varphi\dashv\vdash\neg J_{{}_{0}}\varphi\leftrightarrow 1$.
  • (N): if $\vdash\varphi$ then $\vdash\square\varphi$.
In this Section, by $\vdash$ we will mean $\vdash_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}$.
Lemma 40.
For $\alpha,\beta$ external formulas, the following is a theorem of $\mathsf{PWK}^{\Box}_{\mathrm{e}}$:
  • (BK)
    $\square(\alpha\to\beta)\to(\square\alpha\to\square\beta)$
Proof.
It is the same of Lemma 11. ∎
The Deduction theorem holding for $\mathsf{PWK}_{\mathrm{e}}$ can be actually extented to its modal version.
Theorem 41 (Deduction Theorem).
For the logic $\mathsf{PWK}^{\Box}_{\mathrm{e}}$, it holds that $\Gamma\vdash\varphi$ iff there exist some formulas $\gamma_{1},\dots,\gamma_{n}\in\Gamma$ such that $\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\to\neg J_{{}_{0}}\varphi$.
Proof.
($\Rightarrow$). By induction on the length of the derivation of $\varphi$ from $\Gamma$.
Basis. If $\varphi$ is an axiom ($\vdash\varphi$), then by Lemma 9 we have $\vdash\neg J_{{}_{0}}\varphi$.
Inductive step. We just show the case of the rule (N). Let $\varphi=\Box\psi$, for some $\psi\in Fm$, and $\Gamma\vdash\Box\psi$, and the last deduction rule applied is (N), hence it holds $\vdash\psi$. Therefore we have $\vdash\Box\psi$, hence $\vdash\neg J_{{}_{0}}\Box\psi$, by Lemma 9. By induction hypothesis, there exist some formulas $\gamma_{1},\dots,\gamma_{n}\in\Gamma$ such that $\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\to\neg J_{{}_{0}}\psi$. Observe that (by classical logic, Lemma 9) $\vdash\neg J_{{}_{0}}\Box\psi\to(\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\to\neg J_{{}_{0}}\Box\psi)$, hence, by modus ponens (on external formulas, see Lemma 9), $\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\to\neg J_{{}_{0}}\Box\psi$.
($\Leftarrow$). Suppose $\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\to\neg J_{{}_{0}}\varphi$. By Lemma 9, we have $\Gamma\vdash\neg J_{{}_{0}}\gamma$, $\forall\gamma\in\Gamma$, therefore $\Gamma\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}$, thus applying modus ponens (on external formulas by Lemma 9), we get $\Gamma\vdash\neg J_{{}_{0}}\varphi$. Again, by Lemma 9 we have $\Gamma\vdash\varphi$ (as $\neg J_{{}_{0}}\varphi\vdash\varphi$). ∎
Lemma 42.
In the logic $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ it holds:
  1. $\square(J_{i}\varphi\land J_{k}\psi)\leftrightarrow\square J_{i}\varphi\land\square J_{k}\psi$ for $i,k\in\{0,1,2\}$.
Proof.
The proof is identical to that of Lemma 13. Observe that even though $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ does not possess a full modus ponens, in this proof we are dealing exclusively with external formulas for which full modus ponens holds (see Lemma 9).
4.1. Semantics
The semantics for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ employs the weak three-valued Kripke models of Definition 15. In particular we consider the following subclass:
Definition 43.
A PWK-Kripke model is a weak three-valued Kripke model ${\langle W,R,v\rangle}$ such that $v$ evaluates formulas of the form $\Box\varphi$ according to:
  • (1)
    $v(w,\Box\varphi)=1$ iff $v(w,\varphi)\neq\nicefrac{{1}}{{2}}$ and $v(s,\varphi)\neq 0$ for every $s\in W$ such that $wRs$.
  • (0)
    $v(w,\Box\varphi)=0$ iff $v(w,\varphi)\neq\nicefrac{{1}}{{2}}$ and there exists $s\in W$ such that $wRs$ and $v(s,\varphi)=0$.
  • ($\nicefrac{{1}}{{2}}$)
    $v(w,\Box\varphi)=\nicefrac{{1}}{{2}}$ iff $v(w,\varphi)=\nicefrac{{1}}{{2}}$.
Within this Section by model we intend PWK-Kripke model.
Remark 44.
Observe that, in every model ${\langle W,R,v\rangle}$ with $w\in W$, it holds that $v(w,\diamond\varphi)=1$ iff $v(w,\varphi)\neq\nicefrac{{1}}{{2}}$ and there exists $s\in W$ such that $wRs$ and $v(s,\varphi)=1$, $v(w,\diamond\varphi)=0$ iff $v(w,\varphi)\neq\nicefrac{{1}}{{2}}$ and $v(s,\varphi)\neq 1$ for every $s\in W$ such that $wRs$.
We denote by $\models^{l}_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}$ the local modal PWK external logic on the class of all frames obtained by taking $\{1,\nicefrac{{1}}{{2}}\}$ as designated values, that is:
Definition 45.
$\Gamma\vDash^{l}_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}\varphi$ iff for all models ${\langle W,R,v\rangle}$ and all $w\in W$, if $v(w,\gamma)\neq 0,\forall\gamma\in\Gamma$, then $v(w,\varphi)\neq 0$.
[…]
Definition 46.
A formula $\varphi$ is satisfied (valid) in a model ${\langle W,R,v\rangle}$ if $v(w,\varphi)\neq 0$, for some (all) $w\in W$.
Validity in a frame and in a class of frames follow Definition 19, with the exception that we consider only PWK-Kripke models built on a frame.
4.2. Completeness and Decidability
The notion of consistency has to be changed for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$, because the sublogic $\mathsf{PWK}$ is paraconsistent. Therefore we substiture consistent (and maximally consistent) sets with non-trivial ones.
Remark 47.
A set $\Gamma\subset Fm$ is non-trivial if there is no formula $\varphi\in Fm$, such that $\Gamma\vdash J_{i}\varphi$ and $\Gamma\vdash\neg J_{i}\varphi$, for any $i\in\{0,1,2\}$. It is called trivial otherwise.
Definition 48.
A non-trivial set $\Gamma$ is maximally non-trivial whenever $\Gamma\subset\Gamma^{\prime}$ implies that $\Gamma^{\prime}$ is trivial.
A meaningful formula is the same of Definition 23) for modal Bochvar logic, by just considering that $\vdash$ here refers to $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ (and not to $\mathsf{B}_{\mathrm{e}}^{\Box}$).
The following is the analogous of Lemma 24, for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ (indeed, the first claims coincide).
Lemma 49.
Let $w$ be a maximally non-trivial set of formulas, then:
  1. (1)
    if $\neg\varphi\not\in w$ and all the variables occurring in $\varphi$ are meaningful in $w$, then $\varphi\in w$;
  2. (2)
    if all the variables of $\varphi$ are covered, then $\varphi$ is meaningful in $w$;
  3. (3)
    if $\vdash\varphi$ then $\neg\varphi\not\in w$;
  4. (4)
    if $\vdash\varphi$ and every variable in $\varphi$ is meaningful then $\varphi\in w$.
Proof.
(3). Suppose that $\vdash\varphi$ and, by contradiction, that $\neg\varphi\in w$. By Lemma 9 $\vdash\neg J_{{}_{0}}\varphi$, thus $w\vdash J_{{}_{0}}\varphi$. On the other hand $w\vdash J_{{}_{0}}\varphi$ (by Lemma 9). But this implies that $w$ is trivial (by Remark 47).
(4) follows from the previous. ∎
Lemma 50.
Let $\Gamma$ be a non-trivial set of formulas. If $\Gamma\cup\{\varphi\}$ is trivial then $\Gamma\vdash\neg\varphi$.
Proof.
Let $\Gamma$ be a non-trivial set of formulas. Suppose $\Gamma\cup\{\varphi\}$ is trivial. Therefore $\Gamma\cup\{\varphi\}\vdash 0$; by Theorem 41, there exist formulas $\gamma_{1},\dots,\gamma_{n}\in\Gamma$ s.t. $\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\land\neg J_{{}_{0}}\varphi\to\neg J_{{}_{0}}0$, where the non-triviality of $\Gamma$ assures that $J_{{}_{2}}\varphi$ actually appears in the antecedent. We have $\vdash_{\mathsf{PWK}_{\mathrm{e}}}\neg J_{{}_{0}}0\leftrightarrow 0$ by Lemma 9, hence $\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\land\neg J_{{}_{0}}\varphi\to 0$, therefore $\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\to J_{{}_{0}}\varphi$. Since $\vdash J_{{}_{0}}\varphi\to\neg J_{{}_{0}}\neg\varphi$, by classical logic $\vdash\neg J_{{}_{0}}\gamma_{1}\wedge\dots\wedge\neg J_{{}_{0}}\gamma_{n}\to\neg J_{{}_{0}}\neg\varphi$. Thus, by Theorem 41, $\gamma_{1}\wedge\dots\wedge\gamma_{n}\vdash\neg\varphi$, hence by monotonicity $\Gamma\vdash\neg\varphi$. ∎
Lindenbaum’s lemma for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ has a slightly different form from that of $\mathsf{B}_{\mathrm{e}}^{\Box}$.
Lemma 51 (Lindenbaum’s Lemma for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$).
Let $\Gamma$ be a non-trivial set of formulas such that $\Gamma\not\vdash\varphi$, for some $\varphi\in Fm$, then there exists a maximally non-trivial set of formulas $w$ such that $\Gamma\subseteq w$ and $\neg\varphi\in w$.
Proof.
Suppose $\Gamma\not\vdash\varphi$. Let $\psi_{1},\psi_{2},\psi_{3},\dots$ be an enumeration of the formulas of $\mathsf{PWK}^{\Box}_{\mathrm{e}}$. We define the sets inductively:
$\Gamma_{0}=\Gamma\cup\{\neg\varphi\}.$
$\Gamma_{i+1}=\left\{\begin{aligned} &\Gamma_{i}\cup\{\psi_{i}\},&&\text{if non-trivial}\\ &\Gamma_{i}\cup\{\neg\psi_{i}\},&&\text{if non-trivial}\\ &\Gamma_{i}\cup\{J_{{}_{1}}\psi_{i}\},&&\text{otherwise}\end{aligned}\right.$
$w=\bigcup\limits_{i\in\mathbb{N}}\Gamma_{i}.$
By construction $w$ is maximal, $\Gamma\subseteq w$ and $\neg\varphi\in w$. We prove the non-triviality of $w$ by induction on $n\in\mathbb{N}$. For the base step, since $\Gamma\not\vdash\varphi$, by Lemma 50 $\Gamma_{0}=\Gamma\cup\{\neg\varphi\}$ is non-trivial. For the inductive step, suppose $\Gamma_{i}$ is non-trivial, while both $\Gamma_{i}\cup\{\psi_{i}\}$ and $\Gamma_{i}\cup\{\neg\psi_{i}\}$ are trivial. Therefore $\Gamma_{i+1}=\Gamma_{i}\cup\{J_{{}_{1}}\psi_{i}\}$. By the same lemma the previous facts yield $\Gamma_{i}\vdash\neg\psi_{i}$ and $\Gamma_{i}\vdash\psi_{i}$. By Lemma 9 we obtain $\Gamma_{i}\vdash J_{{}_{2}}\neg\psi_{i}\lor J_{{}_{1}}\neg\psi_{i}$ and $\Gamma_{i}\vdash J_{{}_{2}}\psi_{i}\lor J_{{}_{1}}\psi_{i}$, of which the former can be rewritten by the same lemma as $\Gamma_{i}\vdash J_{{}_{0}}\psi_{i}\lor J_{{}_{1}}\psi_{i}$. Since $J_{{}_{0}}\psi_{i},J_{{}_{1}}\psi_{i},J_{{}_{2}}\psi_{i}$ are pairwise contradictory, by classical logic we conclude $\Gamma_{i}\vdash J_{{}_{1}}\psi_{i}$. Since $\Gamma_{i}$ is taken as non-trivial, this implies $\Gamma_{i}\nvdash\neg J_{{}_{1}}\psi_{i}$. By Lemma 50, we conclude that $\Gamma_{i}\cup\{J_{{}_{1}}\psi_{i}\}=\Gamma_{i+1}$ is non-trivial. ∎
Definition 52.
Let $\mathcal{W}$ be the set of all maximally non-trivial set of formulas. Then the canonical relation $\mathcal{R}\subset\mathcal{W}\times\mathcal{W}$ for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ is defined, for every $w,s\in\mathcal{W}$ as:
$w\mathcal{R}s\text{ iff }\forall\varphi\in Fm\text{ s.t. }\Box\varphi\in w\text{ then }\neg\varphi\notin s.$
Lemma 53 (Existence Lemma).
For every maximally non-trivial set of formulas $w$ such that $\diamond\varphi\in w$ (for some $\varphi\in Fm$) then $\varphi$ is meaningful in $w$ and there exists a maximally non-trivial set $s\in\mathcal{W}$ such that $w\mathcal{R}s$ and $\varphi\in s$.
Proof.
Let $\diamond\varphi\in w$ and consider the set
$s^{-}=\{\psi\;|\;\Box\neg J_{{}_{0}}\psi\in w\}.$
Observe that $s^{-}\neq\emptyset$, in fact by Lemma 9, $\vdash\neg J_{{}_{0}}1$, therefore by (N) $\vdash\Box\neg J_{{}_{0}}1$. Since $w$ is maximally non-trivial, $\Box\neg J_{{}_{0}}1\in w$, hence $1\in s^{-}$. Suppose by contradiction that $s^{-}$ is trivial. Therefore $s^{-}$ derives every formula, in particular $s^{-}\vdash\neg\varphi$. By Theorem 41, there are $\psi_{1},\dots,\psi_{n}\in s^{-}$ s.t. $\vdash\neg J_{{}_{0}}\psi_{1}\land\dots\land\neg J_{{}_{0}}\psi_{n}\to\neg J_{{}_{0}}\neg\varphi$. By (N) ($\vdash\Box(\neg J_{{}_{0}}\psi_{1}\land\dots\land\neg J_{{}_{0}}\psi_{n}\to\neg J_{{}_{0}}\neg\varphi)$, and using (BK) and modus ponens (since the formulas considered here are external) we get ($\vdash\Box(\neg J_{{}_{0}}\psi_{1}\land\dots\land\neg J_{{}_{0}}\psi_{n})\to\Box\neg J_{{}_{0}}\neg\varphi$. Now Lemma 42 can be generalized to $\vdash\Box\neg J_{{}_{0}}\psi_{1}\land\dots\land\Box\neg J_{{}_{0}}\psi_{n}\to\Box(\neg J_{{}_{0}}\psi_{1}\land\dots\land\neg J_{{}_{0}}\psi_{n})$, obtaining, by transitivity, $\vdash\Box\neg J_{{}_{0}}\psi_{1}\land\dots\land\Box\neg J_{{}_{0}}\psi_{n}\to\Box\neg J_{{}_{0}}\neg\varphi$. Observe that $\Box\neg J_{{}_{0}}\psi_{1},\dots,\Box\neg J_{{}_{0}}\psi_{n}\in w$, therefore $\Box\neg J_{{}_{0}}\neg\varphi\in w$ (as $w$ is maximally non-trivial). By (P2) $\Box\neg\varphi\in w$, which can be rewritten as $\neg\diamond\varphi\in w$, contradicting the non-triviality of $w$.
Notice that we have also proved that in particular $s^{-}\nvdash\neg\varphi$. We can now apply Lindembaum’s lemma and extend $s^{-}$ to a maximally non-trivial $s\supseteq s^{-}$ s.t. $\varphi\in s$, hence $\varphi$ is meaningful in $s$. Finally we show that $w\mathcal{R}s$ according to the canonical relation: let for arbitrary $\chi\in Fm,\Box\chi\in w$, then by (P2) $\Box\neg J_{{}_{0}}\chi\in w$, so $\chi\in s^{-}\subseteq s$ and by non-triviality $\neg\chi\notin s$. ∎
We now adapt the definition of canonical model to $\mathsf{PWK}^{\Box}_{\mathrm{e}}$:
Definition 54.
The canonical model for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ is a weak three-valued Kripke model
$\mathcal{M}={\langle\mathcal{W},\mathcal{R},v\rangle}$ where $\mathcal{W}$ is the set of all maximally non-trivial sets of formulas, $\mathcal{R}$ is the canonical relation for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ and $v$ is defined as follows:
  • $v(w,x)=1$ if and only if $x\in w$;
  • $v(w,x)=0$ if and only if $\neg x\in X$;
  • $v(w,x)=\nicefrac{{1}}{{2}}$ if and only if $J_{{}_{1}}x\in w$,
for every $w\in\mathcal{W}$ and propositional variable $x$.
Lemma 55 (Truth Lemma).
Let $\mathcal{M}={\langle\mathcal{W},\mathcal{R},v\rangle}$ be the canonical model. Then, for every formula $\varphi\in Fm$ and every $w\in\mathcal{W}$, the following hold:
  1. (1)
    $v(w,\varphi)=1$ if and only if $\varphi\in w$;
  2. (2)
    $v(w,\varphi)=0$ if and only if $\neg\varphi\in w$;
  3. (3)
    $v(w,\varphi)=\nicefrac{{1}}{{2}}$ if and only if $J_{{}_{1}}\varphi\in w$.
Proof.
By induction on the length of the formula $\varphi$. We just show (1) for the inductive step when $\varphi=\Box\psi$, for some $\psi\in Fm$.
Observe that $v(w,\Box\psi)=1$ iff $\psi$ is meaningful in $w$ (i.e. $v(w,\psi)\neq\nicefrac{{1}}{{2}}$) and $\forall s$ s.t. $w\mathcal{R}s$, $v(s,\psi)\neq 0$, thus, by induction hypothesis, iff $J_{{}_{1}}\psi\not\in w$ and $\neg\psi\notin s$ $\forall s$ s.t. $w\mathcal{R}s$.
()($\Rightarrow$)( ⇒ ) Suppose, by contradiction, that $v(w,\Box\psi)=1$ but $\Box\psi\not\in w$, hence, by maximality of $w$, $J_{{}_{1}}\Box\psi\in w$ or $\neg\Box\psi\in w$. Since $\psi$ is meaningful in $w$, so is $\Box\psi$ (Lemma 49), thus $J_{{}_{1}}\Box\psi\not\in w$. So, $\neg\Box\psi\in w$, i.e. $\diamond\neg\psi\in w$, hence, by Existence Lemma 53, there exists $s^{\prime}\in\mathcal{W}$ such that $w\mathcal{R}s^{\prime}$ such that $\neg\psi\in s^{\prime}$, which implies, by induction hypothesis, that $v(s^{\prime},\psi)=0$, a contradiction.
()($\Leftarrow$)( ⇐ ) The proof is similar to one of the Truth Lemma 32 for $\mathsf{B}_{\mathrm{e}}^{\square}$. Observe that $\vdash_{\mathsf{PWK}_{\mathrm{e}}}\varphi\to J_{{}_{2}}\varphi$ and $\vdash_{\mathsf{PWK}_{\mathrm{e}}}J_{{}_{2}}\varphi\to+\varphi$, moreover maximally non-trivial sets are closed under unrestricted modus ponens. ∎
Theorem 56 (Completeness).
$\Gamma\vdash_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}\varphi$ if and only if $\Gamma\models^{l}_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}\varphi$.
Proof.
()($\Rightarrow$)( ⇒ ) It is easily checked that all the axioms are sound and the rules preserve soundness.
()($\Leftarrow$)( ⇐ ) Suppose $\Gamma\not\vdash\varphi$. Then $\Gamma$ is a non-trivial set of formulas, therefore, by the Lindenbaum Lemma 51, there exist a maximally non-trivial set $s$ such that $\Gamma\subseteq s$ and $\neg\varphi\in s$, hence, by Truth Lemma 32, there exists a canonical countermodel, namely $(\mathcal{W},\mathcal{R},v)$, with $v(s,\gamma)=1$ for all $\gamma\in\Gamma$ and $v(s,\varphi)=0$. ∎
The decidability of $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ is proved similarly to the case for $\mathsf{B}_{\mathrm{e}}^{\square}$. Once again we make use of filtrations. The notion of closure under subformulas is the same given in Definition 34. The only essential change concerns the equivalence relation that induces the partition.
Definition 57.
Let ${\langle W,R,v\rangle}$ be a model and $\Sigma$ be a finite set of formulas closed under subformulas. This set induces an equivalence relation over $W$ such that: $w\equiv_{\Sigma}s$ iff $\forall\varphi\in\Sigma(v(w,\varphi)=1$ iff $v(s,\varphi)\neq 0)$.
Again, when there is no risk of confusion we omit the reference set $\Sigma$ and denote the equivalence class $[w]_{\equiv_{\Sigma}}$ as $[w]$. We can adopt Definition 36 for filtration.
Theorem 58.
Let ${\langle W^{f},R^{f},v^{f}\rangle}$ be a filtration of $M={\langle W,R,v\rangle}$ through $\Sigma$. For all $\varphi\in\Sigma,w\in W$, it holds $v(w,\varphi)\neq 0$ iff $v^{f}([w],\varphi)\neq 0$.
Proof.
By induction on the complexity of $\varphi\in\Sigma$. The Boolean cases and the case for $\varphi=J_{{}_{2}}\psi$ follows Lemma 37.
Let $\varphi=\square\psi$, for some $\psi\in Fm$. When $v(w,\square\psi)=1$ the proof is similar to the case covered in Lemma 37. Suppose $v(w,\square\psi)=\nicefrac{{1}}{{2}}$: this holds iff $v(w,\psi)=\nicefrac{{1}}{{2}}$ iff (by induction hypothesis) $v([w],\psi)=\nicefrac{{1}}{{2}}$ iff $v([w],\square\psi)=\nicefrac{{1}}{{2}}$. ∎
In accordance with the notion of satisfiability in $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ we reobtain the main theorem:
Theorem 59.
If a formula $\varphi$ is satisfiable in a model, it is satisfiable in a finite model.
Proof.
Identical to Theorem 38. ∎
Corollary 60 (Decidability).
The logic $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ is decidable.
5. Extensions of modal weak Kleene logics
The aim of this Section is to axiomatize some extensions of both the modal logics $\mathsf{B}_{\mathrm{e}}^{\Box}$ and $\mathsf{PWK}^{\Box}_{\mathrm{e}}$. In particular, we focus on those extensions whose semantical counterpart is given by reflexive, transitive and/or Euclidean models (clearly, the properties refer all to the relational part of models). To this end, consider the following formulas:
[…]
which substantially consist of “external versions” [^6] The problem with standard modal formulas is the same with Bochvar (non-external) propositional logic, which is well-known as a logic without theorems, since every formula can be evaluated into $\nicefrac{{1}}{{2}}$. Similarly, in the modal context using the internal formulas (T), (4) or (5) would provide an unsound axiomatization. of the standard modal formulas (T), (4) and (5). The formulas introduced above allows to capture some frame properties within $\mathsf{B}_{\mathrm{e}}^{\square}$, as shown by the following.
Proposition 61.
Let $\mathcal{F}=(W,R)$ be a frame. Then
  1. (1)
    $\mathcal{F}\vDash_{\mathsf{B}_{\mathrm{e}}^{\square}}$ Te iff $R$ is reflexive;
  2. (2)
    $\mathcal{F}\vDash_{\mathsf{B}_{\mathrm{e}}^{\square}}$ 4e iff $R$ is transitive;
  3. (3)
    $\mathcal{F}\vDash_{\mathsf{B}_{\mathrm{e}}^{\square}}$ 5e iff $R$ is euclidean.
Proof.
[…]
()($\Leftarrow$)( ⇐ ) It is immediate to check that Te is valid in every reflexive frame.
(2) ()($\Rightarrow$)( ⇒ ) Assume $\mathcal{F}\models$ 4e, let $w\in W$ such that $wRs^{\prime}$ and $s^{\prime}\mathcal{R}t$, for some $s^{\prime},t\in W$. Consider the valuation $v(s,p)=1$ iff $wRs$, for every $s\in W$ and any propositional variable $p$. This implies that $v(s,J_{{}_{2}}p)=1$ for every $wRs$, thus $v(w,\Box J_{{}_{2}}p)=1$ (observing that $v(w,J_{{}_{2}}p)\neq\nicefrac{{1}}{{2}}$. Since $\mathcal{F}\models$ 4e then $v(w,\Box\Box J_{{}_{2}}p)=1$. Since $wRs^{\prime}$, then $v(s^{\prime},\Box J_{{}_{2}}p)=1$, therefore $v(t,J_{{}_{2}}p)=1$ (since $s^{\prime}Rt$), thus $v(t,p)=1$. This implies that $wRt$, i.e. $R$ is transitive as desired.
()($\Leftarrow$)( ⇐ ) It is immediate to check that 4e is valid in every transitive frame.
(3) ()($\Rightarrow$)( ⇒ ) Suppose $\mathcal{F}$ to be non-euclidean, therefore for some $w,s^{\prime},s^{\prime\prime}\in W,wRs^{\prime},wRs^{\prime\prime}$ but ${\langle s^{\prime},s^{\prime\prime}\rangle}\notin R$. Define the valuation $v$ such that for an arbitrary variable $p$, $v(w,p)=v(s^{\prime\prime},p)=1$, while $v(s^{\prime},p)=0$ and for all $t$ s.t. $s^{\prime}Rt,v(t,p)=0$. It follows that $v(w,\diamond J_{{}_{2}}p)=1$ but $v(s^{\prime},\diamond J_{{}_{2}}p)=0$, therefore $v(w,\Box\diamond J_{{}_{2}}p)=0$, hence $v(w,\diamond J_{{}_{2}}p\to\Box\diamond J_{{}_{2}}p)=0$. This countermodel proves $\mathcal{F}\nvDash\text{5}_{e}.$
()($\Leftarrow$)( ⇐ ) It is immediate to check that 5e is valid in every euclidean frame. ∎
In the case of the logic $\mathsf{PWK}^{\Box}_{\mathrm{e}}$, the same frame properties are expressed by the standard modal formulas:
[…]
Proposition 62.
Let $\mathcal{F}=(W,R)$ be a frame. Then
  1. (1)
    $\mathcal{F}\vDash_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}$ (T) iff $R$ is reflexive;
  2. (2)
    $\mathcal{F}\vDash_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}$ (4) iff $R$ is transitive;
  3. (3)
    $\mathcal{F}\vDash_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}$(5) iff $R$ is euclidean.
Proof.
The proofs runs similarly as Proposition 61. ∎
[…]
Theorem 63.
The relation $\mathcal{R}$ of the canonical models [^7] The canonical model for a certain extension $\mathsf{B}_{\mathrm{e}}^{\square}$x of $\mathsf{B}_{\mathrm{e}}^{\square}$ differs from Definition 31 only for the set of worlds, which now consist not of all maximal consistent sets (w.r.t. $\mathsf{B}_{\mathrm{e}}^{\square}$), but only of the maximal consistent theories of $\mathsf{B}_{\mathrm{e}}^{\square}$x. The canonical model for $\mathsf{PWK}^{\Box}_{\mathrm{e}}$x is adapted from Definition 54 in a similar fashion. for the following logics have the properties:
[…]
Proof.
It follows from Propositions 61 and 62. ∎
That the accessibility relation of the canonical model has the desired properties is enough to obtain the following completeness results:
Corollary 64.
The following hold:
  • $\mathsf{B}_{\mathrm{e}}^{\square}\emph{T}_{e}$ is complete with respect to the class of reflexive frames;
  • $\mathsf{B}_{\mathrm{e}}^{\square}4_{e}$ is complete with respect to the class of transitive frames;
  • $\mathsf{B}_{\mathrm{e}}^{\square}5_{e}$ is complete with respect to the class of euclidean frames;
  • $\mathsf{B}_{\mathrm{e}}^{\square}\emph{S}4_{e}$ is complete with respect to the class of reflexive and transitive frames;
  • $\mathsf{B}_{\mathrm{e}}^{\square}\emph{S}5_{e}$ is complete with respect to the class of frames whose relation is an equivalence.
Corollary 65.
The following hold:
[…]
  • $\mathsf{PWK}^{\Box}_{\mathrm{e}}4$ is complete with respect to the class of transitive frames;
  • $\mathsf{PWK}^{\Box}_{\mathrm{e}}5$ is complete with respect to the class of euclidean frames;
[…]
Notice that depending on the choice of the basic logic we obtain a different notion of completeness: in the case of $\mathsf{B}_{\mathrm{e}}^{\square}$ the completeness is w.r.t. the logical consequence relation $\vDash_{\mathsf{B}_{\mathrm{e}}^{\square}}$, in the case of $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ the relation is $\vDash_{\mathsf{PWK}^{\Box}_{\mathrm{e}}}$.
The decidability of $\mathsf{B}_{\mathrm{e}}^{\square}$ and $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ established by Corollaries 39 and 60 immediately follows for their (finitely axiomatizable) axiomatic extensions.
Theorem 66.
[…]
Proof.
We give the proof for $\mathsf{B}_{\mathrm{e}}^{\square}$e, the others cases employs the same strategy. Using the same definitions of set closed under subformulas and filtration from Definitions 34 and 36, we have that Lemma 37 still holds. Suppose that $\varphi$ is satisfiable in a reflexive model $M$, therefore by Proposition 61 Te is valid in $M$. Let $\Gamma=\{\text{T}_{e},\varphi\}$, $\Sigma$ its closure under subformulas, and consider the filtration $M^{\Sigma}$ of $M$ through $\Sigma$. By Lemma 37, Te is valid in $M^{\Sigma}$, hence by Proposition 61 $M^{\Sigma}$ is reflexive. Now we repeat Theorem 38 using the filtration $M^{\Sigma}$ to prove that if $\varphi$ is satisfiable in reflexive model, it is satisfiable in a finite reflexive model of cardinality at most $2^{card(\Sigma)}$. We have the desired finite model property, which, together with the completeness theorem stated in Corollary 64, gives the decisability of $\mathsf{B}_{\mathrm{e}}^{\square}$e. ∎
6. Conclusions and future work
In this work we have studied two modal expansions of the external weak Kleene logics $\mathsf{B}_{\mathrm{e}}$ and $\mathsf{PWK}_{\mathrm{e}}$. The modalization yields two different operators $\Box$, which, without further inquiry, can be intended as generic necessary (in the alethic sense) operators. The distinction between the two is motivated by the difference in designated values within the propositional bases. Moreover, the logics $\mathsf{B}_{\mathrm{e}}$ and $\mathsf{PWK}_{\mathrm{e}}$ have been first provided with new axiomatizations, which differ from their ones (by Finn-Grigolia and Segerberg, respectively) since they are obtained thanks to the recent results that state the algebraizability of $\mathsf{B}_{\mathrm{e}}$ ([6]) and $\mathsf{PWK}_{\mathrm{e}}$ (see Section 7).
The introduction of modalities to $\mathsf{B}_{\mathrm{e}}$ and $\mathsf{PWK}_{\mathrm{e}}$ adds further expressive power to a language already capable of expressing a (sort of) truth predicate ($J_{{}_{2}}$), allowing to speak about the “truth” of a formula. The interplay between $J_{{}_{2}}$ and $\Box$ makes these logics able to express some interactions between the notion of necessity and truth. Let us observe that, at first glance, external connectives could look very much like the “statability” operators introduced by Correia in [19] for the (internal) modal version of PWK. However, the substantial difference is that statability operators are actually modal operators, while external connectives work at the propositional level. Nevertheless, the language of our logics is rich enough to allow the defition of (at least) one of them, namely Correia’s statability operator $\mathcal{S}$ can be defined as $\mathcal{S}(\varphi)\coloneqq\Box+\varphi$.
The logics $\mathsf{B}_{\mathrm{e}}^{\square}$ and $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ have been presented syntactically via Hilbert-style systems and provided with a possible worlds semantics in terms of three-valued Kripke models. Completeness and decidability of the axiomatic calculi have been established, the former via Henkin-style proofs, the latter by the filtration technique.
As it is customary in modal logic, some standard axiomatic extensions of the logics have been presented. In the case of $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ this is done proving that the well-known formulas T,4, and 5 correspond to their usual properties on the accessibility relation. On $\mathsf{B}_{\mathrm{e}}$ we introduced the counterparts of those formulas in the external language and prove their correspondence with frame properties. The idea of considering these extensions goes in the direction of exploring epistemic logics that could be useful to formalize notions such as knowledge and ignorance in a non-classical context (see e.g. [33], [6]). A more general question about if and how standard modal formulas defining frame properties can be transferred (see [12]) to $\mathsf{B}_{\mathrm{e}}^{\square}$ and $\mathsf{PWK}^{\Box}_{\mathrm{e}}$ obtaining classes of frames characterized by the same properties is an interesting topic yet to be explored.
In the paper the algebraizability of $\mathsf{PWK}_{\mathrm{e}}$ has been proved as well. This seemingly side result is the starting point for a further work on the algebraic semantics for $\mathsf{B}_{\mathrm{e}}^{\square}$ and $\mathsf{PWK}^{\Box}_{\mathrm{e}}$. Together with the algebraizability of $\mathsf{B}_{\mathrm{e}}$, the result of the current paper completes the strong correspondence between external weak Kleene logics and their algebraic counterpart, the class of Bochvar algebras.
[…]
Finally, let us emphasize that we deliberately chose to work with weak Kleene logics as they are “traditionally” intended, namely the consequence relations defined via either truth or non-falsity preservation. A different choice to be explored in the future consists in modalizing logics of “mixed” nature, such as those where a non-false consequence can follow from true (only) premises (see e.g. [16]), a possibility that is briefly discussed at the end of [19].
7. Appendix: Algebraizability of $\mathsf{PWK}_{\mathrm{e}}$
Recall from Definition 2 that $\mathsf{PWK}_{e}$ is the logic induced by the matrix $\langle\textbf{WK}^{e},\{1,\nicefrac{{1}}{{2}}\}\rangle$. A Bochvar algebra is an algebra ${\mathbf{A}}=\langle A,\lor,\land,\neg,J_{{}_{2}},0,1\rangle$ of type $\langle 2,2,1,1,0,0\rangle$ satisfying the following identities and quasi-identities. [^9] Bochvar algebras were originally introduced by Finn and Grigolia [21] in an extended language including $J_{{}_{1}}$ and $J_{{}_{0}}$. These two operations are term-definable in the language $\langle A,\lor,\land,\neg,J_{{}_{2}},0,1\rangle$, as already explained in Section 2. Besides, we are using here the much shorter equivalent axiomatization introduced in [10, Theorem 7].
  1. (1)
    $\varphi\vee\varphi\thickapprox\varphi$;
  2. (2)
    $\varphi\lor\psi\thickapprox\psi\lor\varphi$;
  3. (3)
    $(\varphi\lor\psi)\lor\delta\thickapprox\varphi\lor(\psi\lor\delta)$;
  4. (4)
    $\varphi\land(\psi\lor\delta)\thickapprox(\varphi\land\psi)\lor(\varphi\land\delta)$;
  5. (5)
    ¬($\neg(\neg\varphi)\thickapprox\varphi$;
  6. (6)
    $\neg 1\thickapprox 0$;
  7. (7)
    $\neg(\varphi\lor\psi)\thickapprox\neg\varphi\land\neg\psi$;
  8. (8)
    $0\vee\varphi\thickapprox\varphi$;
  9. (9)
    $J_{{}_{0}}J_{{}_{2}}\varphi\thickapprox\neg J_{{}_{2}}\varphi$;
  10. (10)
    $J_{{}_{2}}\varphi\thickapprox\neg(J_{{}_{0}}\varphi\vee J_{{}_{1}}\varphi)$;
  11. (11)
    $J_{{}_{2}}\varphi\vee\neg J_{{}_{2}}\varphi\thickapprox 1$;
  12. (12)
    $J_{{}_{2}}(\varphi\lor\psi)\thickapprox(J_{{}_{2}}\varphi\land J_{{}_{2}}\psi)\vee(J_{{}_{2}}\varphi\land J_{{}_{2}}\neg\psi)\vee(J_{{}_{2}}\neg\varphi\land J_{{}_{2}}\psi)$;
  13. (13)
    $J_{{}_{0}}\varphi\thickapprox J_{{}_{0}}\psi\;\&\;J_{{}_{2}}\varphi\thickapprox J_{{}_{2}}\psi\;\Rightarrow\;\varphi\thickapprox\psi$.
Bochvar algebras form a proper quasi-variety, which we denote by $\mathsf{BCA}$, and that is generated by $\textbf{WK}^{e}$. It plays the role of equivalent algebraic semantics of both Bochvar external logic [6, Theorem 35] and PWK external logic, as we show in Theorem 67. Recall that a logic $\vdash_{\mathsf{L}}$ is algebraizable ([26, Definition 3.11, Proposition 3.12]) with respect to a certain class of algebras $\mathsf{K}$ if there exist two maps $\tau\colon Fm\to\mathcal{P}(Eq)$, $\rho\colon Eq\to\mathcal{P}(Fm)$ from formulas to sets of equations and from equations to sets of formulas such that:
[…]
and
[…]
As a notational convention, in the following proof we will indicate by $Hom(\mathbf{Fm},{\mathbf{WK}}^{e})$ the set of homomorphisms from the formula algebra (in the language of $\mathsf{BCA}$) in ${\mathbf{WK}}^{e}$.
Theorem 67.
$\mathsf{PWK}_{\mathrm{e}}$ is algebraizable w.r.t. $\mathsf{BCA}$ with transformers $\tau(\varphi):=\{\neg J_{{}_{0}}\varphi\approx 1\}$ and $\rho(\epsilon\approx\delta):=\{\epsilon\equiv\delta\}$.
Proof.
In our case (ALG1) and (ALG4) translate into:
  • (ALG1)
    $\Gamma\vdash_{\mathsf{PWK}_{e}}\varphi\Leftrightarrow\tau[\Gamma]\vDash_{{}_{\mathsf{BCA}}}\tau(\varphi)$,
  • (ALG4)
    $\varepsilon\approx\delta\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}\vDash_{{}_{\mathsf{BCA}}}\tau(\rho(\varepsilon\approx\delta))$.
Moreover, since $\mathsf{BCA}$ is the quasi-variety generated by ${\mathbf{WK}}^{e}$, the above claims amount to the following:
  • (ALG1)
    $\Gamma\vdash_{\mathsf{PWK}_{e}}\varphi\Leftrightarrow\tau[\Gamma]\vDash_{{\mathbf{WK}}^{e}}\tau(\varphi)$,
  • (ALG4)
    $\varepsilon\approx\delta\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}\vDash_{{\mathbf{WK}}^{e}}\tau(\rho(\varepsilon\approx\delta))$.
(ALG1) ($\Rightarrow$) Suppose $\Gamma\vdash_{\mathsf{PWK}_{e}}\varphi$. Take $h\in Hom(\mathbf{Fm},{\mathbf{WK}}^{e})$ s.t. $h(\neg J_{{}_{0}}\gamma)=h(1)=1$ for every $\gamma\in\Gamma$, which implies $\neg J_{{}_{0}}h(\gamma)=1$, i.e. $h(\gamma)\neq 0$. Since $\Gamma\vdash_{\mathsf{PWK}_{e}}\varphi$, by Definition 2 it holds $h(\varphi)\neq 0$, hence $h(\neg J_{{}_{0}}\varphi)=1=h(1)$. Thus, we have shown that $\tau[\Gamma]\vDash_{{\mathbf{WK}}^{e}}\tau(\varphi)$.
($\Leftarrow$) Suppose $\tau[\Gamma]\vDash_{{\mathbf{WK}}^{e}}\tau(\varphi)$ and let $h\in Hom(\mathbf{Fm},{\mathbf{WK}}^{e})$ be s.t. $h(\gamma)\neq 0$ for every $\gamma\in\Gamma$. Therefore, by the hypothesis, $h(\neg J_{{}_{0}}\varphi)=h(1)=1$, which implies $h(\varphi)\neq 0$, giving the desired conclusion.
(ALG4) ($\Rightarrow$) Consider an arbitrary identity $\varepsilon\approx\delta$ in the language of $\mathsf{BCA}$. A simple calculation shows that $\tau(\rho(\varepsilon\approx\delta))$ is $\neg J_{{}_{0}}(\varepsilon\equiv\delta)\approx 1$. Let $h\in Hom(\mathbf{Fm},{\mathbf{WK}}^{e})$ be s.t. $h(\varepsilon)=h(\delta)$, therefore $h(\varepsilon\equiv\delta)=1$. This implies $h(\neg J_{{}_{0}}(\varepsilon\equiv\delta))=1=h(1)$, so we conclude $\varepsilon\approx\delta\models_{{\mathbf{WK}}^{e}}\tau(\rho(\varepsilon\approx\delta))$.
()($\Leftarrow$)( ⇐ ) Suppose $h(\neg J_{{}_{0}}(\varepsilon\equiv\delta))=\neg J_{{}_{0}}(h(\varepsilon)\equiv h(\delta))=1$, for $h\in Hom(\mathbf{Fm},{\mathbf{WK}}^{e})$. Now $J_{{}_{0}}(h(\varepsilon)\equiv h(\delta))=0$ implies $(J_{{}_{2}}h(\varepsilon)\leftrightarrow J_{{}_{2}}h(\delta))\land(J_{{}_{0}}h(\varepsilon)\leftrightarrow J_{{}_{0}}h(\delta))\in\{1,\nicefrac{{1}}{{2}}\}$, but since $(J_{{}_{2}}h(\varepsilon)\leftrightarrow J_{{}_{2}}h(\delta))\land(J_{{}_{0}}h(\varepsilon)\leftrightarrow J_{{}_{0}}h(\delta))$ is an external formula it must be that $(J_{{}_{2}}h(\varepsilon)\leftrightarrow J_{{}_{2}}h(\delta))\land(J_{{}_{0}}h(\varepsilon)\leftrightarrow J_{{}_{0}}h(\delta))=1$. The fact that $J_{{}_{2}}h(\varepsilon)\leftrightarrow J_{{}_{2}}h(\delta)=1$ and $J_{{}_{0}}h(\varepsilon)\leftrightarrow J_{{}_{0}}h(\delta)=1$ implies $J_{{}_{2}}h(\varepsilon)=J_{{}_{2}}h(\delta)$ and $J_{{}_{0}}h(\varepsilon)=J_{{}_{0}}h(\delta)$. Applying the quasi-equation (13), we conclude $h(\varepsilon)=h(\delta)$.
Acknowledgments. We thank two anonymous referees for their detailed comments and observations, which allowed to improve the quality of the paper.
S. Bonzio acknowledges the support by the Italian Ministry of Education, University and Research through the PRIN 2022 project DeKLA (“Developing Kleene Logics and their Applications”, project code: 2022SM4XC8) and the PRIN Pnrr project “Quantum Models for Logic, Computation and Natural Processes (Qm4Np)” (cod. P2022A52CR). He also acknowledges the Fondazione di Sardegna for the support received by the projects GOACT (grant number F75F21001210007) and MAPS (grant number F73C23001550007), the University of Cagliari for the support by the StartUp project “GraphNet”, and also the support by the MOSAIC project (H2020-MSCA-RISE-2020 Project 101007627). Finally, he gratefully acknowledges also the support of the INDAM GNSAGA (Gruppo Nazionale per le Strutture Algebriche, Geometriche e loro Applicazioni).
N. Zamperlin acknowledges the support of the MSCA-RISE Programme – Plexus: Philosophical, Logical, and Experimental Routes to Substructurality.